### v1

# Tri-resonant leptogenesis from modular symmetry neutrino models

Abstract
In this paper, we have studied the consequences of some representative modular symmetry neutrino models for tri-resonant leptogenesis (which is realized by having three nearly degenerate right-handed neutrinos). To be specific, we have considered modular ${\rm A}_{4}$, ${\rm S}_{4}$ and ${\rm A}_{5}$ symmetry models that have a right-handed neutrino mass matrix $M_{\rm R}$ as shown in Eq. (8) which gives three degenerate right-handed neutrino masses and consequently prohibits the leptogenesis mechanism to work. For these models, we have considered two minimal ways to generate the desired small mass splittings among the three right-handed neutrinos: one scenario is to modify $M_{\rm R}$ to a form as shown in Eq. (10), while another scenario is to consider the renormalization-group corrections for the right-handed neutrino masses. In the former scenario, for the considered models, the observed value of $\eta_{\rm B}$ can always be successfully reproduced for appropriate values of $\mu$. But in the latter scenario not all the considered models can successfully reproduce the observed value of $\eta_{\rm B}$.
1 Introduction
As we know, the discovery of neutrino oscillations reveals that neutrinos are massive and their flavor eigenstates $\nu_{\alpha}$ (for $\alpha=e,\mu,\tau$) are certain superpositions of their mass eigenstates $\nu_{i}$ (for $i=1,2,3$) possessing definite masses $m_{i}$: $\nu_{\alpha}=\sum_{i}U_{\alpha i}\nu_{i}$ with $U_{\alpha i}$ being the $\alpha i$-th element of the neutrino mixing matrix $U$ [1]. In the standard form, $U$ is expressed in terms of three mixing angles $\theta_{ij}$ (for $ij=12,13,23$), one Dirac CP phase $\delta$ and two Majorana CP phases $\alpha_{21}$ and $\alpha_{31}$ as
$U=\left(\begin{matrix}c_{12}c_{13}&s_{12}c_{13}&s_{13}e^{-{\rm i}\delta}\cr-s_{12}c_{23}-c_{12}s_{23}s_{13}e^{{\rm i}\delta}&c_{12}c_{23}-s_{12}s_{23}s_{13}e^{{\rm i}\delta}&s_{23}c_{13}\cr s_{12}s_{23}-c_{12}c_{23}s_{13}e^{{\rm i}\delta}&-c_{12}s_{23}-s_{12}c_{23}s_{13}e^{{\rm i}\delta}&c_{23}c_{13}\end{matrix}\right)\left(\begin{matrix}1&&\cr&e^{{\rm i}\alpha_{21}}&\cr&&e^{{\rm i}\alpha_{31}}\end{matrix}\right)\;,$ (1)
where the abbreviations $c_{ij}=\cos\theta_{ij}$ and $s_{ij}=\sin\theta_{ij}$ have been employed.
Nowadays, the neutrino mixing angles and neutrino mass squared differences $\Delta m^{2}_{ij}\equiv m^{2}_{i}-m^{2}_{j}$ have been measured to good degrees of accuracy, and there is also a preliminary result for $\delta$ (but with a large uncertainty). Several research groups have performed global analyses of the existing neutrino oscillation data to extract the values of these parameters [2, 3]. For reference, the results in Ref. [2] are reproduced in Table 1 here. Note that the sign of $\Delta m^{2}_{31}$ remains undetermined, thereby allowing for two possible neutrino mass orderings: the normal ordering (NO) $m_{1}<m_{2}<m_{3}$ and inverted ordering (IO) $m_{3}<m_{1}<m_{2}$. However, neutrino oscillations are completely insensitive to the absolute neutrino mass scale and the Majorana CP phases. Their values can only be inferred from certain non-oscillatory experiments such as the neutrinoless double beta decay experiments [4]. But so far there has not been any lower bound on the lightest neutrino mass, nor any constraint on the Majorana CP phases.
For the neutrino mixing angles, Table 1 shows that $\theta_{12}$ and $\theta_{23}$ are close to some special values (i.e., $\sin^{2}\theta_{12}\sim 1/3$ and $\sin^{2}\theta_{23}\sim 1/2$). But $\theta_{13}$ is relatively small. In fact, before its value was measured, $\theta_{13}$ had been widely expected to be vanishingly small. For the ideal case of $\sin^{2}\theta_{12}=1/3$, $\sin^{2}\theta_{23}=1/2$ and $\theta_{13}=0$, the neutrino mixing matrix takes a very simple form as
$U_{\rm TBM}=\displaystyle\frac{1}{\sqrt{6}}\left(\begin{array}[]{ccc}2&\sqrt{2}&0\cr 1&-\sqrt{2}&-\sqrt{3}\cr 1&-\sqrt{2}&\sqrt{3}\cr\end{array}\right)\;,$ (5)
which is referred to as the tribimaximal (TBM) mixing [5]. This particular mixing has inspired intensive model-building studies with the employment of certain discrete non-Abelian symmetries (e.g., ${\rm A}_{4}$ and ${\rm S}_{4}$) [6].
Noteworthy, a shortcoming of the conventional flavor-symmetry models is that many flavon fields need to be introduced in order to achieve a desired breaking pattern of the flavor symmetry, and the scalar potential of these flavon fields needs a complicated symmetric shaping. Moreover, higher-dimensional operators with more flavon insertions and unknown coefficients are often present, affecting the predictive power of the models. In recent years, the idea of modular invariance which provides substantial simplifications to these problems has been proposed and become increasingly popular [7, 8]. In the modular-invariance approach, the flavour symmetry is broken by the VEV of a single scalar field—the so-called modulus, while the flavon fields are not necessary. And the Yukawa couplings and fermion mass terms are certain special functions of the modulus called modular forms that have particular properties under the modular group transformations. Moreover, all higher-dimensional operators are unambiguously determined in the limit of unbroken supersymmetry. As a result, flavor models with modular invariance depend on fewer parameters, enhancing their predictive power.
As for the neutrino masses, one of the most popular and natural ways of generating them is the type-I seesaw model [9] in which three heavy right-handed neutrinos $N_{I}$ ($I=1,2,3$) are introduced into the Standard Model (SM). First of all, $N_{I}$ can constitute the Yukawa coupling operators together with the lepton doublets $L_{\alpha}$ and the Higgs doublet $H$: $Y_{I\alpha}\overline{N_{I}}HL_{\alpha}$ with $Y_{I\alpha}$ being the $I\alpha$-th element of the Yukawa coupling matrix $Y$. These operators will generate the Dirac neutrino mass terms $(M_{\rm D})_{I\alpha}=Y_{I\alpha}v$ [here $(M_{\rm D})_{I\alpha}$ is the $I\alpha$-th element of the Dirac neutrino mass matrix $M_{\rm D}$] after the neutral component of $H$ acquires a nonzero vacuum expectation value (VEV) $v=174$ GeV. In addition, $N_{I}$ themselves can also have the Majorana mass terms $\overline{N_{I}}(M_{\rm R})_{IJ}N^{c}_{J}$ [here $(M_{\rm R})_{IJ}$ are the $IJ$-th element of the right-handed neutrino mass matrix $M_{\rm R}$]. Then, under the seesaw condition $M_{\rm R}\gg M_{\rm D}$, integrating the right-handed neutrinos out yields an effective Majorana mass matrix for three light neutrinos as
$M_{\nu}=-M^{T}_{\rm D}M^{-1}_{\rm R}M_{\rm D}\;.$ (6)
Here we would like to point out that throughout the paper the fermion mass matrices are written in the right-left convention which is more conveniently used in the supersymmetric setup as applicable to our study.
Normal Ordering Inverted Ordering
bfp ±$\pm 1\sigma$ $3\sigma$ range bfp ±$\pm 1\sigma$ $3\sigma$ range
$\sin^{2}\theta_{12}$ $0.303_{-0.011}^{+0.012}$ $0.270\to 0.341$ $0.303_{-0.011}^{+0.012}$ $0.270\to 0.341$
$\sin^{2}\theta_{23}$ $0.572_{-0.023}^{+0.018}$ $0.406\to 0.620$ $0.578_{-0.021}^{+0.016}$ $0.412\to 0.623$
$\sin^{2}\theta_{13}$ $0.02203_{-0.00059}^{+0.00056}$ $0.02029\to 0.02391$ $0.02219_{-0.00057}^{+0.00060}$ $0.02047\to 0.02396$
$\delta/\pi$ $1.09_{-0.14}^{+0.23}$ $0.96\to 1.33$ $1.59_{-0.18}^{+0.15}$ $1.41\to 1.74$
$\Delta m^{2}_{21}/(10^{-5}~{}{\rm eV}^{2})$ $7.41_{-0.20}^{+0.21}$ $6.82\to 8.03$ $7.41_{-0.20}^{+0.21}$ $6.82\to 8.03$
$\Delta m^{2}_{3\ell}/(10^{-3}~{}{\rm eV}^{2})$ $2.511_{-0.027}^{+0.028}$ $2.428\to 2.597$ $-2.498_{-0.025}^{+0.032}$ $-2.581\to-2.408$
[…]
Remarkably, the seesaw model also provides an attractive explanation (which is known as the leptogenesis mechanism [10, 11]) for the baryon-antibaryon asymmetry of the Universe [12]
$Y_{\rm B}\equiv\frac{n_{\rm B}-n_{\rm\bar{B}}}{s}\simeq(8.69\pm 0.04)\times 10^{-11}\;,$ (7)
where $n_{\rm B}$ ($n_{\rm\bar{B}}$) denotes the baryon (antibaryon) number density and $s$ the entropy density. The leptogenesis mechanism works in a way as follows: a lepton-antilepton asymmetry $Y_{\rm L}\equiv(n_{\rm L}-n_{\rm\bar{L}})/s$ is firstly generated from the out-of-equilibrium and CP-violating decays of the right-handed neutrinos and then partly converted into the baryon-antibaryon asymmetry via the sphaleron processes: $Y_{\rm B}\simeq-cY_{\rm L}$ with $c=28/79$ or $8/23$ in the SM or MSSM (Minimal Supersymmetric Standard Model) framework. As is known, in the scenario that the right-handed neutrino masses are hierarchial there exists a lower bound about $10^{9}$ GeV for the right-handed neutrino mass scale in order to successfully reproduce the observed value of $Y_{\rm B}$ [13]. Obviously, such a high right-handed neutrino mass scale has no chance to be directly accessed by forseeable collider experiments. Fortunately, in the resonant leptogenesis scenario which is realized for nearly degenerate right-handed neutrinos [14], a successful leptogenesis can be realized even with TeV-scale right-handed neutrinos which have the potential to be directly accessed by running or upcoming experiments [15].
However, for some typical modular symmetry neutrino models (see the models provided in sections 3-5), the right-handed neutrino mass matrix takes a form as
$M_{\rm R}=\begin{pmatrix}M_{0}&0&0\\ 0&0&M_{0}\\ 0&M_{0}&0\end{pmatrix}\;.$ (8)
It can be diagonalized by the following unitary matrix
$U_{\rm R}=\begin{pmatrix}0&1&0\\ \frac{\rm i}{\sqrt{2}}&0&\frac{1}{\sqrt{2}}\\ -\frac{\rm i}{\sqrt{2}}&0&\frac{1}{\sqrt{2}}\end{pmatrix}\;,$ (9)
to give three degenerate right-handed neutrino masses (i.e., $M_{0}$). In this case, unfortunately, the leptogenesis mechanism is prohibited to work. But just as a coin has two sides, this kind of models can serve as a unique basis for realizations of the tri-resonant leptogenesis scenario: if the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent, then the tri-resonant leptogenesis scenario will be naturally realized For a realization of resonant leptogenesis in the minimal seesaw model with a modular symmetry, see Ref. [16]. .
In this paper, we consider two minimal ways to generate the desired small mass splittings among the three right-handed neutrinos: one way is to modify $M_{\rm R}$ to a form as
$M_{\rm R}=\begin{pmatrix}M_{0}&0&0\\ 0&\mu&M_{0}\\ 0&M_{0}&\mu\end{pmatrix}\;.$ (10)
with $\mu\ll M_{0}$. Such a modification of $M_{\rm R}$ has the merit that it only introduces one additional parameter and it still can be diagonalized by the unitary matrix $U_{\rm R}$ in Eq. (9) (i.e., it does not alter the original neutrino mixing pattern). But the three right-handed neutrino masses are modified to $M_{0}-\mu$, $M_{0}$ and $M_{0}+\mu$, respectively. Another way is to consider the renormalization-group corrections for the right-handed neutrino masses [17]: if the energy scale $\Lambda_{\rm FS}$ of the flavor-symmetry physics that shapes the special textures of neutrino mass matrices is much higher than the right-handed neutrino mass scale $M_{0}$ where leptogenesis takes place (such a scenario is very likely given that the idea of modular flavor symmetry is just inspired from the string theory [7]), then the renormalization group evolution (RGE) effect between the scales of $\Lambda_{\rm FS}$ and $M_{0}$ may induce some mass splittings among the three right-handed neutrinos (see section 2.2). Such a correction of $M_{\rm R}$ has the following two merits: 1) this effect is spontaneous, provided that there is a considerable gap between the flavor-symmetry scale and the right-handed neutrino mass scale; 2) this effect is minimal, in the sense that it does not need to introduce additional parameters.
The remaining parts of this paper are organized as follows. In section 2 we will first give some basic formulas about resonant leptogenesis and RGE effects that will be used in our study. Then, we perform the study for three representative modular symmetry models which contain three degenerate right-handed neutrinos. We first consider the modular ${\rm A}_{4}$ symmetry model in section 3, then the modular ${\rm S}_{4}$ symmetry model in section 4, and finally the modular ${\rm A}_{5}$ symmetry model in section 5. In section 6, we summarize our main results.
2 Resonant leptogenesis and RGE effects
2.1 Some basics about resonant leptogenesis
As is known, if the temperature $T$ of the Universe at which leptogenesis takes place (approximately the right-handed neutrino mass scale) is below $10^{9}$ GeV, all the three lepton flavors have become distinguishable from one another and the lepton asymmetries stored in them should be tracked separately [18]. Such a three-flavor regime applies to the scenario studied in this paper (i.e., low-scale tri-resonant leptogenesis). In this case, the final lepton asymmetry can be approximately expressed as
$Y_{\rm L}\simeq r(\varepsilon_{e}\kappa_{e}+\varepsilon_{\mu}\kappa_{\mu}+\varepsilon_{\tau}\kappa_{\tau})\;.$ (11)
Here the factor $r\simeq 4\times 10^{-3}$ measures the ratio of the equilibrium number density of right-handed neutrinos to the entropy density. And $\varepsilon_{e}$ is a sum over three right-handed neutrinos (i.e., $\varepsilon_{e}=\varepsilon_{1e}+\varepsilon_{2e}+\varepsilon_{3e}$) of the following CP asymmetries:
$\varepsilon_{I\alpha}\equiv\frac{\Gamma\left(N_{I}\to L_{\alpha}+H\right)-\Gamma\left(N_{I}\to\overline{L}_{\alpha}+\overline{H}\right)}{\sum_{\alpha}\Gamma\left(N_{I}\to L_{\alpha}+H\right)+\Gamma\left(N_{I}\to\overline{L}_{\alpha}+\overline{H}\right)}\;,$ (12)
whose explicit expressions will be given below. Finally, $\kappa_{\alpha}$ are the efficiency factors (which are smaller than 1) due to the washout effects. Their concrete values can be calculated by numerically solving relevant Boltzmann equations, whose explicit expressions will also be given below.
To be explicit, in the basis where both the right-handed neutrino matrix and charged lepton mass matrix are diagonal, the CP asymmetries $\varepsilon_{I\alpha}$ defined in Eq. (12) can be expressed as [14]
$\varepsilon_{I\alpha}=\frac{\left|(\bar{Y}_{+})_{I\alpha}\right|^{2}-\left|(\bar{Y}_{-})_{I\alpha}\right|^{2}}{\left(\bar{Y}_{+}\bar{Y}^{\dagger}_{+}\right)_{II}+\left(\bar{Y}_{-}\bar{Y}^{\dagger}_{-}\right)_{II}}\;.$ (13)
Here $(\bar{Y}_{+})_{I\alpha}$ are the effective neutrino Yukawa couplings in the case of there existing three nearly degenerate right-handed neutrinos. They can be expressed as
$\left(\bar{Y}_{+}\right)_{I\alpha}=Y_{I\alpha}-{\rm i}V_{I\alpha}+{\rm i}\sum_{J,K=1}^{3}\left|\epsilon_{IJK}\right|Y_{J\alpha}$
×$\times\frac{M_{I}\left(\mathcal{M}_{IIJ}+\mathcal{M}_{JJI}\right)+{\rm i}R_{IK}\left[\mathcal{M}_{IKJ}\left(\mathcal{M}_{IIK}+\mathcal{M}_{KKI}\right)+\mathcal{M}_{JJK}\left(\mathcal{M}_{IKI}+\mathcal{M}_{KIK}\right)\right]}{M^{2}_{I}-M^{2}_{J}-2{\rm i}M^{2}_{I}A_{JJ}-2{\rm i}{\rm Im}(R_{IK})\left[M^{2}_{I}\left|A_{JK}\right|^{2}+M_{J}M_{K}{\rm Re}(A^{2}_{JK})\right]}\;,$ (14)
where $\epsilon_{IJK}$ is the anti-symmetric Levi-Civita symbol, $\mathcal{M}_{IJK}\equiv M_{I}A_{JK}$ and
[…]
$R_{IJ}=\frac{M^{2}_{I}}{M^{2}_{I}-M^{2}_{J}-2{\rm i}M^{2}_{I}A_{JJ}}\;,$ (15)
with $f(x)=\sqrt{x}[1-(1+x)\ln(1+1/x)]$ being the loop function. And the corresponding CP-conjugate effective Yukawa couplings $\left(\bar{Y}_{-}\right)_{I\alpha}$ can be obtained by simply making the replacement $Y_{I\alpha}\to Y^{*}_{I\alpha}$ in the above expressions. Furthermore, in terms of these effective Yukawa couplings, the decay widths of right-handed neutrinos can be expressed as
$\Gamma\left(N_{I}\to L_{\alpha}+H\right)=\frac{M_{I}}{16\pi}\left|\left(\bar{Y}_{+}\right)_{I\alpha}\right|^{2}\;,\hskip 28.45274pt\Gamma\left(N_{I}\to\overline{L}_{\alpha}+\overline{H}\right)=\frac{M_{I}}{16\pi}\left|\left(\bar{Y}_{-}\right)_{I\alpha}\right|^{2}\;,$ (16)
and their branching ratios are given by
$B_{I\alpha}=\frac{\left|(\bar{Y}_{+})_{I\alpha}\right|^{2}+\left|(\bar{Y}_{-})_{I\alpha}\right|^{2}}{\left(\bar{Y}_{+}\bar{Y}^{\dagger}_{+}\right)_{II}+\left(\bar{Y}_{-}\bar{Y}^{\dagger}_{-}\right)_{II}}\;.$ (17)
In order to generate a nonzero lepton asymmetry in the Universe, one needs not only the above CP asymmetries, but also a departure from thermal equilibrium. In the latter processes, the evolutions of the number densities of the right-handed neutrinos $n_{N_{I}}$ and flavored lepton asymmetries $n_{L_{\alpha}}$ (which are normalized by the photon number density $n_{\gamma}$ to $\eta_{N_{I}}\equiv n_{N_{I}}/n_{\gamma}$ and $\eta_{L_{\alpha}}\equiv n_{L_{\alpha}}/n_{\gamma}$, respectively) can be tracked by the following Boltzmann equations [14]
$\frac{{\rm d}\eta_{N_{I}}}{{\rm d}z}$ $=$ $\frac{z}{H(z=1)}\Bigg{[}\left(1-\frac{\eta_{N_{I}}}{\eta_{N_{I}}^{\mathrm{eq}}}\right)\sum_{\alpha=e,\mu,\tau}\left(\Gamma^{D(I\alpha)}+\Gamma_{\text{Yukawa }}^{S(I\alpha)}+\Gamma_{\text{Gauge }}^{S(I\alpha)}\right)$
$-\frac{2}{3}\sum_{\alpha=e,\mu,\tau}\eta_{L_{\alpha}}\varepsilon_{I\alpha}\left(\hat{\Gamma}^{D(I\alpha)}+\hat{\Gamma}_{\text{Yukawa }}^{S(I\alpha)}+\hat{\Gamma}_{\text{Gauge }}^{S(I\alpha)}\right)\Bigg{]}\;,$
$\frac{{\rm d}\eta_{L_{\alpha}}}{{\rm d}z}$ $=$ $\frac{z}{H(z=1)}\Bigg{\{}\sum^{3}_{I=1}\varepsilon_{I\alpha}\left(\frac{\eta_{N_{I}}}{\eta_{N_{I}}^{\mathrm{eq}}}-1\right)\sum_{\beta=e,\mu,\tau}\left(\Gamma^{D(I\beta)}+\Gamma_{\text{Yukawa }}^{S(I\beta)}+\Gamma_{\text{Gauge }}^{S(I\beta)}\right)$ (18)
$-\frac{2}{3}\eta_{L_{\alpha}}\Bigg{[}\sum^{3}_{I=1}B_{I\alpha}\left(\tilde{\Gamma}^{D(I\alpha)}+\tilde{\Gamma}_{\text{Yukawa }}^{S(I\alpha)}+\tilde{\Gamma}_{\text{Gauge }}^{S(I\alpha)}+\Gamma_{\text{Yukawa }}^{W(I\alpha)}+\Gamma_{\text{Gauge }}^{W(I\alpha)}\right)$
$\hskip 42.67912pt+\sum_{\beta=e,\mu,\tau}\left(\Gamma_{\text{Yukawa }}^{\Delta L=2(\alpha\beta)}+\Gamma_{\text{Yukawa }}^{\Delta L=0(\alpha\beta)}\right)\Bigg{]}$
$-\frac{2}{3}\sum_{\beta=e,\mu,\tau}\eta_{L_{\beta}}\Bigg{[}\sum^{3}_{I=1}\varepsilon_{I\alpha}\varepsilon_{I\beta}\left(\Gamma_{\text{Yukawa }}^{W(I\beta)}+\Gamma_{\text{Gauge }}^{W(I\beta)}+\Gamma_{\text{Yukawa }}^{\Delta L=2(\alpha\beta)}-\Gamma_{\text{Yukawa }}^{\Delta L=0(\alpha\beta)}\right)\Bigg{]}\Bigg{\}}\;,$
where $z=M_{1}/T$ has been defined, and
$H\left(z\right)=\sqrt{\frac{4\pi^{3}g_{*}}{45}}\frac{M^{2}_{1}}{z^{2}M_{\mathrm{Pl}}}\;,$ (19)
is the Hubble parameter with $M_{\mathrm{Pl}}=1.221\times 10^{19}\mathrm{GeV}$ and $g_{*}\simeq 106.75$ or 228.75 (in the SM and MSSM frameworks, respectively) being the number of relativistic degrees of freedom in the early Universe. And the analytical expressions of the $\Gamma$ terms on the right-hand side of Eq. (18) can be found in Ref. [14]. Note that at the present epoch one has the relation $s\simeq 7.04n_{\gamma}$ for the entropy and photon number densities, and correspondingly $Y_{\rm L}\simeq\eta_{\rm L}/7.04$ (with $\eta_{\rm L}=\eta_{L_{e}}+\eta_{L_{\mu}}+\eta_{L_{\tau}}$).
2.2 Some basics about RGE effects
As mentioned in the introduction section, the flavor-symmetry models which contain three degenerate right-handed neutrinos serve as a unique basis for realizations of the tri-resonant leptogenesis scenario after the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent. And the renormalization group evolution effects from high to low energies can play the role of generating the desired right-handed neutrino mass splitings. In the literature, most of the flavor-symmetry models have been formulated in the MSSM framework for the following reason: in order to break the flavor symmetry in a proper way, one needs to introduce some flavon fields which transform as multiplets of the flavor symmetry and develop particular VEV alignments; and the most popular and perhaps natural approach to derive the desired flavon VEV alignments is provided by the so-called F-term alignment mechanism which is realized in a supersymmetric setup [6]. In particular, all of the modular flavor-symmetry models have been formulated in the MSSM framework since the modular flavor symmetry needs supersymmetry to preserve the holomorphicity of the modular form [7]. For these reasons, we will perform our study in the MSSM framework.
In the MSSM framework, the one-loop RGE evolution of the right-handed neutrino mass matrix and the Dirac neutrino Yukawa matrix $Y$ in the basis where $M_{\rm R}$ is diagonal are given by [19]
$4\pi^{2}\frac{{\rm d}M_{I}}{{\rm d}t}=\left(YY^{\dagger}\right)_{II}M_{I}\;,$
$16\pi^{2}\frac{{\rm d}Y}{{\rm d}t}=-T^{*}Y-Y\left[\alpha-\left(3Y^{\dagger}Y+Y^{\dagger}_{l}Y_{l}\right)\right]\;,$ (20)
where $t=\ln\left(M_{0}/\mu\right)$ has been defined, with $M_{0}$ and $\mu$ denoting the right-handed neutrino mass scale and renormalization scale, respectively. $\alpha=3g^{2}_{2}+\frac{3}{5}g^{2}_{1}-{\rm Tr}\left(3Y^{\dagger}_{\rm U}Y_{\rm U}+Y^{\dagger}Y\right)$, $g_{2}$ and $g_{1}$ are the $\rm SU\left(2\right)_{\rm L}$ and $\rm U\left(1\right)_{\rm Y}$ gauge coupling constants, and $Y_{{\rm U},l}$ are the Yukawa matrices for up-type quarks and charged leptons. $T$ is a $3\times 3$ anti-Hermitian matrix and its entries are given by
$T_{II}=0\;,$
$T_{IJ}=-T^{*}_{JI}=\frac{M_{J}+M_{I}}{M_{J}-M_{I}}{\rm Re}[(YY^{\dagger})_{IJ}]+{\rm i}\frac{M_{J}-M_{I}}{M_{J}+M_{I}}{\rm Im}[(YY^{\dagger})_{IJ}]\;,\qquad{\rm for}\;\;I\neq J\;.$ (21)
$T_{IJ}$ is singular at the scale where $M_{J}=M_{I}$ holds unless ${\rm Re}[(YY^{\dagger})_{IJ}]=0$. To achieve this condition, we perform an additional rotation in the above basis:
$M^{\prime}_{\rm R}=O^{\dagger}M_{\rm R}O^{*}=M_{\rm R}\;,\qquad Y^{\prime}=O^{\dagger}Y\;,$ (22)
where $O$ is a real orthogonal matrix which can be parameterized by three angles $\theta_{1,2,3}$. This rotation does not affect the right-handed neutrino mass matrix due to their exact mass degeneracy. $YY^{\dagger}$ becomes
$Y^{\prime}Y^{\prime\dagger}=O^{\dagger}YY^{\dagger}O\;.$ (23)
Apparently, the three condition ${\rm Re}[(YY^{\dagger})_{IJ}]=0$ can be always satisfied by choosing suitable vuales of $\theta_{1,2,3}$. The RGE evolution of the right-handed neutrino mass matrix in the new basis is
$4\pi^{2}\frac{{\rm d}M_{I}}{{\rm d}t}=\left(Y^{\prime}Y^{\prime\dagger}\right)_{II}M_{I}\;.$ (24)
For illustration, taking account of the RGE effects from the flavor-symmetry scale $\Lambda_{\rm FS}$ to the right-handed neutrino mass scale $M_{0}$, one can obtain the following relation for the right-handed neutrino masses
$M_{I}\left(M_{0}\right)\simeq M_{I}\left(\Lambda_{\rm FS}\right)\left[1-\frac{1}{4\pi^{2}}\left(Y^{\prime}Y^{\prime\dagger}\right)_{II}\ln\left(\frac{\Lambda_{\rm FS}}{M_{0}}\right)\right]\;,$ (25)
yielding the following mass splittings for the right-handed neutrinos
$M_{I}\left(M_{0}\right)-M_{J}\left(M_{0}\right)\simeq\frac{M_{I}\left(\Lambda_{\rm FS}\right)}{4\pi^{2}}\left[\left(Y^{\prime}Y^{\prime\dagger}\right)_{JJ}-\left(Y^{\prime}Y^{\prime\dagger}\right)_{II}\right]\ln\left(\frac{\Lambda_{\rm FS}}{M_{0}}\right)\;.$ (26)
Such mass splittings among the right-handed neutrinos are comparable to their decay widths
$\Gamma_{I}=\frac{\left(YY^{\dagger}\right)_{II}}{8\pi}M_{I}\left(M_{0}\right)\;,$ (27)
fulfilling the condition for the resonant leptogenesis to proceed. Note that in obtaining Eq. (25), as a rough estimate, we have only considered the diagonal-term [i.e., $(YY^{\dagger})_{II}$-related] contributions to the right-handed neutrino masses. But the off-diagonal-term [i.e., $(YY^{\dagger})_{IJ}$-related] contributions may be equally important in the case of the right-handed neutrino masses being degenerate, so in the following calculations we will also consider them.
3 Study for modular ${\rm A}_{4}$ symmetry model
In Ref. [20], the authors have presented a comprehensive analysis of neutrino masses and mixing in theories with modular ${\rm A}_{4}$ symmetry. ${\rm A}_{4}$ is the symmetry group of the even permutations of 4 objects, and it has three singlet representations ${\bf 1}$, ${\bf 1^{\prime}}$, ${\bf 1^{\prime\prime}}$ and one three-dimensional representation ${\bf 3}$. Note that ${\rm A}_{4}$ is the smallest non-Abelian finite group which admits a three-dimensional representation (a desired feature for describing the flavor mixing of three generation fermions). In this model, both the three left-handed lepton doublets and three right handed neutrinos constitute the ${\bf 3}$ representation of ${\rm A}_{4}$. In the case that the weights of the lepton doublets and right-handed neutrinos are $(k_{L},k_{N^{c}})=(2,0)$, the modular invariance constrains the Dirac neutrino mass matrix to the following form
$M_{D}=v_{u}\begin{pmatrix}2g_{1}Y_{1}&\left(-g_{1}+g_{2}\right)Y_{3}&\left(-g_{1}-g_{2}\right)Y_{2}\\ \left(-g_{1}-g_{2}\right)Y_{3}&2g_{1}Y_{2}&\left(-g_{1}+g_{2}\right)Y_{1}\\ \left(-g_{1}+g_{2}\right)Y_{2}&\left(-g_{1}-g_{2}\right)Y_{1}&2g_{1}Y_{3}\end{pmatrix}\;,$ (28)
while the right-handed neutrino mass matrix takes the same form as in Eq. (8). On the other hand, in the case that three right-handed charged leptons are successively assigned to the singlet representations ${\bf 1}$, ${\bf 1^{\prime\prime}}$ and ${\bf 1^{\prime}}$ of $A_{4}$, the charged-lepton mass matrix takes a form as
$M_{l}=v_{d}\begin{pmatrix}\alpha Y_{1}&\alpha Y_{3}&\alpha Y_{2}\\ \beta Y_{2}&\beta Y_{1}&\beta Y_{3}\\ \gamma Y_{3}&\gamma Y_{2}&\gamma Y_{1}\end{pmatrix}.$ (29)
In the above two equations, $v_{u}$ and $v_{d}$ are respectively the VEVs of the up-type and down-type Higgs fields, $g_{1}$, $g_{2}$, $\alpha$, $\beta$ and $\gamma$ are simply dimensionless parameters, while $Y_{i}$ are modular forms of level 3 and they can be expressed in terms of the Dedekind eta-function
$\eta(\tau)\equiv q^{1/24}\prod_{n=1}^{\infty}\left(1-q^{n}\right)\hskip 14.22636pt{\rm with}\hskip 14.22636ptq=e^{{\rm i}2\pi\tau}\;,$ (30)
and its derivative (with ${\rm d}\eta/{\rm d}\tau$ being denoted as $\eta^{\prime}$) as
$Y_{1}(\tau)=\frac{\rm i}{2\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\frac{\eta^{\prime}(\tau/3+1/3)}{\eta(\tau/3+1/3)}+\frac{\eta^{\prime}(\tau/3+2/3)}{\eta(\tau/3+2/3)}-\frac{27\eta^{\prime}(3\tau)}{\eta(3\tau)}\right]\;,$ (31)
$Y_{2}(\tau)=\frac{-i}{\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\omega^{2}\frac{\eta^{\prime}(\tau/3+1/3)}{\eta(\tau/3+1/3)}+\omega\frac{\eta^{\prime}(\tau/3+2/3)}{\eta(\tau/3+2/3)}\right]\;,$ (32)
$Y_{3}(\tau)=\frac{-i}{\pi}\left[\frac{\eta^{\prime}(\tau/3)}{\eta(\tau/3)}+\omega\frac{\eta^{\prime}(\tau/3+1/3)}{\eta(\tau/3+1/3)}+\omega^{2}\frac{\eta^{\prime}(\tau/3+2/3)}{\eta(\tau/3+2/3)}\right]\;,$ (33)
with $\omega=e^{{\rm i}2\pi/3}$. One can see that this model totally contains 8 real parameters: ${\rm Re}(\tau)$, ${\rm Im}(\tau)$, $\beta/\alpha$, $\gamma/\alpha$, $|g_{2}/g_{1}|$, ${\rm arg}(g_{2}/g_{1})$, $\alpha v_{d}$ and $g^{2}_{1}v^{2}_{u}/M_{0}$. The three parameters $\alpha v_{d}$, $\beta/\alpha$ and $\gamma/\alpha$ can be fixed by fitting with the values of three charged lepton masses. The remaining parameters serve to describe the three neutrino masses, three neutrino mixing angles and three CP phases present in the neutrino mixing matrix. The detailed numerical calculation shows that this model can be consistent with the neutrino oscillation experimental results in both the NO and IO cases. And the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 2.
NO IO
Best-fit Allowed regions Best-fit Allowed regions
${\rm Re}(\tau)$ 0.0129 $[0,0.431]$ 0.096 $[0,0.102]$
${\rm Im}(\tau)$ 1.824 $[0.91,1.16]\cup[1.31,1.86]$ 0.987 $[0.98,1.049]\cup[1.052,1.109]$
$\beta/\alpha$ 205.720 $[192.39,215]\cup[3054.254093.49]$ 79.472 $[59.68,86.37]\cup[892.67,1446.02]$
$\gamma/\alpha$ 3612.07 $[192.4,215]\cup[3066.47,4092.98]$ 1232.57 $[60.97,86.34]\cup[870.16,1443.84]$
$|g_{2}/g_{1}|$ 2.410 $[2.398,2.71]\cup[2.95,3.86]$ 2.093 $[1.038,2.453]$
$\arg{(g_{2}/g_{1})}$ 6.267 $[0,0.49]\cup[6.23,2\pi]$ 4.715 $[1.33,1.83]\cup[4.29,5]$
$\alpha v_{d}$/MeV 0.5179 1.167
$(g^{2}_{1}v_{u}^{2}/M_{0})$/eV 0.0111 0.004
$m_{e}/m_{\mu}$ 0.0048 $[0.0046,0.0050]$ 0.0048 $[0.0046,0.0050]$
$m_{\mu}/m_{\tau}$ 0.0564 $[0.0520,0.0610]$ 0.0565 $[0.0520,0.0610]$
$\sin^{2}\theta_{12}$ 0.3096 $[0.2750,0.3500]$ 0.3100 $[0.2750,0.3500]$
$\sin^{2}\theta_{13}$ 0.0226 $[0.02045,0.02439]$ 0.02264 $[0.02068,0.02463]$
$\sin^{2}\theta_{23}$ 0.4638 $[0.4180,0.4676]$ 0.584 $[0.423,0.629]$
$\delta/\pi$ 1.486 $[0.19,0.77]\cup[1.23,1.84]$ 1.458 $[0.068,1.933]$
$\alpha_{21}/\pi$ 0.068 $[0,0.23]\cup[1.77,2]$ 0.138 $[0,0.19]\cup[1.8,2]$
$\alpha_{31}/\pi$ 0.948 $[0.683,1.311]$ 0.997 $[0,2]$
$m_{1}$/eV 0.0430 $[0.0225,0.0478]$ 0.0494 $[0.0464,0.0526]$
$m_{2}$/eV 0.0439 $[0.0241,0.0485]$ 0.0501 $[0.0472,0.0533]$
$m_{3}$/eV 0.0661 $[0.0524,0.0716]$ 0.0013 $[0.0007,0.0015]$
$\chi^{2}_{\mathrm{min}}$ 30.72 $10^{-7}$
Table 2: For the modular ${\rm A}_{4}$ symmetry model given in section 3, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in the NO and IO cases. These results are taken from Tables 6 and 10 in Ref. [20].
We first consider the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10). In order to facilitate the leptogenesis calculation, we go to the basis where both the right-handed neutrino mass matrix and charged lepton mass matrix are diagonal with the help of $U_{\rm R}$ in Eq. (9) (which serves to diagonalize the the right-handed neutrino mass matrix) and another unitary matrix $U_{\rm L}$ which serves to diagonalize the charged lepton mass matrix. At the same time, the neutrino Yukawa coupling matrix becomes $Y^{\prime}=U^{\dagger}_{\rm R}YU_{\rm L}$ which is relevant for the leptogenesis calculations. With help of the Eq. (13) and the fitted values of the input parameters as given in Table 2, we numerically calculate the CP asymmetries in the decays of $N_{I}$ as functions of $2\mu/\Gamma_{I}$. The results are shown in Figure 1. Since $\mu$ is a free parameter, the resonance conditions $2\mu/\Gamma_{I}\simeq 1$ which enhance the CP asymmetries to their maximally allowed values can always be fulfilled. This will be helpful for us to reproduce the observed value of $\eta_{\rm B}$. Figure 2 has further shown the ratios of various collision terms in the Boltzmann equations to the Hubble constant as functions of $z=M_{0}/T$ with $M_{0}=10^{3}$ GeV and $\mu=10^{-8}$ GeV as benchmark values (see below for the reason why we have taken $\mu=10^{-8}$ GeV as a benchmark value). We see that the scattering collision term $\Gamma^{S}_{\rm Gauge}$ is comparable to $\Gamma^{S}_{\rm Yukawa}$, and they provide the dominant sources of the scattering collision terms in the temperature range $T>M_{0}$. On the other hand, in the temperature range $T<M_{0}$, the dominant term is $\Gamma^{D}_{\rm Yukawa}$ that describes the decay process. Then, the numerical scan finds the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ as shown in Figure 3. Considering that people are more interested in low-scale leptogenesis (which has the potential to be directly accessed by running or upcoming experiments), our scan over $M_{0}$ is performed in the range from $10^{3}$ GeV to $10^{6}$ GeV. The results show that for each value of $M_{0}$ there exist two values of $\mu$ that accommodate a successful leptogenesis, and the ratio of $\mu$ to $M_{0}$ is about $10^{-10}$ or $10^{-14}$ ($10^{-9}$ or $10^{-21}$) in the NO (IO) cases. For the purpose of illustration, by taking $M_{0}=10^{3}$ GeV as a benchmark value, in Figure 4 we have shown the evolutions of $\eta_{N}$ and $\eta_{B}$ as functions of $z$. These results are obtained by taking the initial conditions to be $\eta^{\rm ini}_{N}=\eta^{\rm ini}_{B}=0$. The results show that, for the benchmark value $M_{0}=10^{3}$ GeV, the observed value of $\eta_{\rm B}$ can be successfully reproduced for $\mu\simeq 10^{-8.4}$ GeV ($10^{-5.9}$ GeV) in the NO (IO) case.
Figure 1: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), CP asymmetries in the decays of $N_{I}$ as functions of the ratio $2\mu/\Gamma_{I}$ in the NO (left panel) and IO (right panel) cases. Here $\varepsilon_{\rm tot}=\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}$ has been defined.
Figure 2: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the various collision terms in the Boltzmann equations as functions of $z=M_{0}/T$ in the NO (left panel) and IO (right panel) cases. These results are obtained by taking $M_{0}=10^{3}$ GeV and $\mu=10^{-8}$ GeV as benchmark values.
Figure 3: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ in the NO (left panel) and IO (right panel) cases.
Figure 4: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the evolutions of $\eta_{N}$ and $\eta_{B}$ as functions of $z$ for the benchmark value $M_{0}=10^{3}$ GeV and some benchmark values of $\mu$ in the NO (left panel) and IO (right panel) cases. Here the horizontal grey dashed line shows the observed value of $\eta_{\rm B}$ (same as below) and the vertical dotted line represents the decoupling temperature $T_{*}\simeq 131.7$ GeV of sphalerons [21].
Then, we consider the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects. In this regard, Figure 5 has shown the RGE-induced mass splittings of the three right-handed neutrinos. In obtaining these results, we have fixed the flavor-symmetry scale $\Lambda_{\rm FS}$ at the grand unification (GUT) scale $10^{16}$ GeV. The results show that the ratios $\Delta M_{IJ}/M_{0}$ (for $\Delta M_{IJ}\equiv M_{I}-M_{J}$) are linearly proportional to $M_{0}$. This is because $\Delta M_{IJ}/M_{0}$ are proportional to $YY^{\dagger}$ [see Eq. (26)], which in turn is proportional to $M_{0}$ because of the seesaw relationship [see Eq. (6)]. Figure 6 further shows the CP asymmetries in the decays of $N_{I}$ as functions of $M_{0}$. In the NO case, for $M_{0}<10^{11}$ GeV, the CP asymmetries are small and nearly independent of $M_{0}$. But for $M_{0}>10^{11}$ GeV, $\varepsilon_{3}$ increases rapidly with $M_{0}$, while $\varepsilon_{1}$ and $\varepsilon_{2}$ also do so after a sign reversal from being positive to negative. In the IO case, the CP asymmetries keep positive and change very slowly as $M_{0}$ increases. Finally, Figure 7 has shown the generated $\eta_{B}$ as functions of $M_{0}$. As one can see, the observed value of $\eta_{B}$ can be successfully reproduced for $M_{0}\simeq 6\times 10^{12}$ GeV ($4\times 10^{12}$ GeV or $6\times 10^{12}$ GeV) in the NO (IO) case. These results can be understood with the help of the results in Figure 6.
Figure 5: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, $\Delta M_{IJ}/M_{0}$ (for $\Delta M_{IJ}\equiv M_{I}-M_{J}$) as functions of $M_{0}$ in the NO (left panel) and IO (right panel) cases.
Figure 6: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the CP asymmetries in the decays of $N_{I}$ as functions of $M_{0}$ in the NO (left panel) and IO (right panel) cases.
Figure 7: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the generated $\eta_{\rm B}$ as functions of $M_{0}$ in the NO (in the red color) and IO (in the blue color) cases.
4 Study for modular ${\rm S}_{4}$ symmetry model
In Ref. [22], the authors have presented a comprehensive analysis of neutrino masses and mixing in theories with modular ${\rm S}_{4}$ symmetry. In the case that the three right-handed neutrinos have a weight $k_{N^{c}}=0$ and constitute the ${\bf 3}$ or ${\bf 3^{\prime}}$ representation of ${\rm S}_{4}$, the right-handed neutrino mass matrix will take a same form as in Eq. (8). When the three left-handed lepton doublets have a weight $k_{L}=2$ and constitute the same three-dimensional representation of ${\rm S}_{4}$ as the three right-handed neutrinos, the Dirac neutrino mass matrix will take a form as
$M_{\rm D}=gv_{u}\left[\begin{pmatrix}0&Y_{1}&Y_{2}\\ Y_{1}&Y_{2}&0\\ Y_{2}&0&Y_{1}\end{pmatrix}+\frac{g^{\prime}}{g}\begin{pmatrix}0&Y_{5}&-Y_{4}\\ -Y_{5}&0&Y_{3}\\ Y_{4}&-Y_{3}&0\end{pmatrix}\right]\;.$ (34)
If the three left-handed lepton doublets constitute a three-dimensional representation of ${\rm S}_{4}$ different from the three right-handed neutrinos, then the Dirac neutrino mass matrix will take a form as
$M_{\rm D}=gv_{u}\left[\begin{pmatrix}0&-Y_{1}&Y_{2}\\ -Y_{1}&Y_{2}&0\\ Y_{2}&0&-Y_{1}\end{pmatrix}+\frac{g^{\prime}}{g}\begin{pmatrix}2Y_{3}&-Y_{5}&-Y_{4}\\ -Y_{5}&2Y_{4}&-Y_{3}\\ -Y_{4}&-Y_{3}&2Y_{5}\end{pmatrix}\right]\;.$ (35)
On the other hand, for the minimal (in terms of weights) possibility, the charged-lepton mass matrix takes a form as
$M_{l}=v_{d}\begin{pmatrix}\alpha Y_{3}&\alpha Y_{5}&\alpha Y_{4}\\ \beta\left(Y_{1}Y_{4}-Y_{2}Y_{5}\right)&\beta\left(Y_{1}Y_{3}-Y_{2}Y_{4}\right)&\beta\left(Y_{1}Y_{5}-Y_{2}Y_{3}\right)\\ \gamma\left(Y_{1}Y_{4}+Y_{2}Y_{5}\right)&\gamma\left(Y_{1}Y_{3}+Y_{2}Y_{4}\right)&\gamma\left(Y_{1}Y_{5}+Y_{2}Y_{3}\right)\\ \end{pmatrix}\;.$ (36)
In the above three equations, $g$, $g^{\prime}$, $\alpha$, $\beta$ and $\gamma$ are simply dimensionless parameters, while $Y_{i}$ are modular forms of weight 2 and level 4 whose explicit expressions have been formulated in Ref. [23].
One can see that this model also contains 8 real parameters in total: $\alpha v_{d}$, $\beta/\alpha$, $\gamma/\alpha$, $g^{2}v^{2}_{u}/M_{0}$, $|g^{\prime}/g|$, ${\rm arg}(g^{\prime}/g)$, ${\rm Re}(\tau)$ and ${\rm Im}(\tau)$. The numerical calculation shows that this model can be consistent with the neutrino oscillation experimental results. For $M_{\rm D}$ in Eq. (34), the model can be consistent with the neutrino oscillation experimental results only in the NO case, and the numerical search finds two pairs of distinct local minima of $\Delta\chi^{2}$ corresponding to two pairs of distinct values of $\tau$. The two minima in each pair lead to opposite values of the Dirac and Majorana phases, but the same values of all other observables. The cases belonging to the first pair are denoted as Case A, while those belonging to the second pair are denoted as Case B. For these two cases, the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 3 [22]. For $M_{\rm D}$ in Eq. (35), the model can be consistent with the neutrino oscillation experimental results in both the NO and IO cases. In the IO case, the numerical search finds two pairs of distinct local minima of $\Delta\chi^{2}$ corresponding to two pairs of distinct values of $\tau$, and the cases belonging to these two pairs are respectively denoted as Case C and Case D. In the NO case, the numerical search only finds one pair of local minima of $\Delta\chi^{2}$ corresponding to one pair of distinct values of $\tau$, and the cases belonging to this pair are denoted as Case E. For these three cases, the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 4 [22].
Case A (NO) Case B (NO)
Best-fit Allowed regions Best-fit Allowed regions
${\rm Re}(\tau)$$\pm 0.1045$$\pm(0.09378-0.1128)$ $\mp 0.109$ $\mp(0.103-0.1197)$
${\rm Im}(\tau)$ 1.01 $1.004-1.018$ 1.005 $0.9988-1.008$
$\beta/\alpha$ 9.465 $7.693-12.39$ 0.03306 $0.02529-0.04074$
$\gamma/\alpha$ 0.002205 $0.001941-0.002472$ 0.0001307 $0.0000982-0.0001663$
${\rm Re}(g^{\prime}/g)$ 0.233 $-0.02544-0.4417$ 0.4097 $0.3241-0.5989$
${\rm Im}(g^{\prime}/g)$$\pm 0.4924$$\pm(-0.6046-0.5751)$ $\mp 0.5745$ $\mp(0.5436-0.5944)$
$\alpha v_{d}$ [MeV] 53.19 893.2
$g^{2}v_{u}^{2}/M_{0}$ [eV] 0.00467 0.004014
$m_{e}/m_{\mu}$ 0.004802 $0.00422-0.005383$ 0.004802 $0.004211-0.005384$
$m_{\mu}/m_{\tau}$ 0.0565 $0.04317-0.06961$ 0.05649 $0.04318-0.06962$
$\sin^{2}\theta_{12}$ 0.305 $0.2656-0.3449$ 0.305 $0.2662-0.3455$
$\sin^{2}\theta_{13}$ 0.02125 $0.01912-0.02383$ 0.02125 $0.01914-0.02383$
$\sin^{2}\theta_{23}$ 0.551 $0.4838-0.5999$ 0.551 $0.4322-0.601$
$m_{1}$ [eV] 0.01746 $0.01185-0.02143$ 0.02074 $0.01918-0.02428$
$m_{2}$ [eV] 0.01945 $0.01473-0.02307$ 0.02244 $0.02101-0.02574$
$m_{3}$ [eV] 0.05288 $0.05075-0.05452$ 0.05406 $0.05314-0.05577$
$\delta/\pi$$\pm 1.314$$\pm(1.249-1.961)$$\pm 1.919$$\pm(1.882-1.977)$
$\alpha_{21}/\pi$$\pm 0.302$$\pm(0.2748-0.3708)$$\pm 1.704$$\pm(1.681-1.722)$
$\alpha_{31}/\pi$$\pm 0.8716$$\pm(0.7973-1.635)$$\pm 1.539$$\pm(1.484-1.618)$
$N\sigma\equiv\sqrt{\Delta\chi^{2}}$ 0.02005 0.02435
Table 3: For the model given in section 4 with $M_{\rm D}$ in Eq. (34), in the phenomenologically viable NO case, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in Case A and Case B. These results are taken from Tables 5a-5b in Ref. [22].
Case C (IO) Case D (IO) Case E (NO)
Best-fit Allowed regions Best-fit Allowed regions Best-fit Allowed regions
${\rm Re}(\tau)$ $\mp 0.1435$ $\mp(0.1222-0.168)$$\pm 0.179$$\pm(0.1589-0.199)$ $\mp 0.4996$ $\mp(0.48-0.5084)$
${\rm Im}(\tau)$ 1.523 $1.088-1.594$ 1.397 $1.236-1.529$ 1.309 $1.246-1.385$
$\beta/\alpha$ 17.82 $9.32-23.66$ 15.35 $10.79-21.09$ 0.000243 $0.0002004-0.0002864$
$\gamma/\alpha$ 0.003243 $0.00227-0.003733$ 0.002924 $0.002443-0.003459$ 0.03335 $0.02799-0.03926$
${\rm Re}(g^{\prime}/g)$ $-0.8714$ $-(0.7956-1.148)$ $-1.32$ $-(1.131-1.447)$ $-0.06454$ $-(0.01697-0.1215)$
${\rm Im}(g^{\prime}/g)$ $\mp 2.094$ $\mp(1.409-2.182)$$\pm 1.733$$\pm(1.306-2.017)$ $\mp 0.569$ $\mp(0.4572-0.6564)$
$\alpha v_{d}$ [MeV] 71.26 68.42 1125
$g^{2}v_{u}^{2}/M_{0}$ [eV] 0.004087 0.00447 0.0087
$m_{e}/m_{\mu}$ 0.004797 $0.004215-0.005378$ 0.004786 $0.004221-0.005386$ 0.004797 $0.004393-0.005197$
$m_{\mu}/m_{\tau}$ 0.05655 $0.04348-0.0698$ 0.0554 $0.04343-0.06968$ 0.05626 $0.04741-0.0654$
$\sin^{2}\theta_{12}$ 0.303 $0.2657-0.3436$ 0.3031 $0.2657-0.3436$ 0.311 $0.2895-0.3375$
$\sin^{2}\theta_{13}$ 0.02175 $0.01957-0.0242$ 0.02184 $0.01954-0.0242$ 0.02185 $0.02041-0.02351$
$\sin^{2}\theta_{23}$ 0.5571 $0.4551-0.6026$ 0.5577 $0.5482-0.6013$ 0.4469 $0.43-0.4614$
$m_{1}$ [eV] 0.0513 $0.04882-0.05207$ 0.05122 $0.05023-0.05212$ 0.01774 $0.01703-0.01837$
$m_{2}$ [eV] 0.05201 $0.04958-0.05274$ 0.05193 $0.05098-0.05279$ 0.0197 $0.01906-0.02025$
$m_{3}$ [eV] 0.01512 $0.00316-0.0163$ 0.01495 $0.01223-0.01649$ 0.05299 $0.05251-0.05346$
$\delta/\pi$$\pm 1.098$$\pm(0.98-1.289)$$\pm 1.384$$\pm(1.271-1.437)$$\pm 1.601$$\pm(1.287-1.828)$
$\alpha_{21}/\pi$$\pm 1.241$$\pm(1.113-1.758)$$\pm 1.343$$\pm(1.171-1.479)$$\pm 1.093$$\pm(0.8593-1.178)$
$\alpha_{31}/\pi$$\pm 0.2487$$\pm(0.069-0.346)$$\pm 0.806$$\pm(0.448-1.149)$$\pm 0.7363$$\pm(0.3334-0.9643)$
$N\sigma\equiv\sqrt{\Delta\chi^{2}}$ 0.0357 0.3811 2.147
Table 4: For the model given in section 4 with $M_{\rm D}$ in Eq. (35), in their respective phenomenologically viable neutrino mass ordering cases, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in Case C, Case D and Case E. These results are taken from Tables 5c-5e in Ref. [22].
Now, in a similar way as in section 3, we study the consequences of these five cases for tri-resonant leptogenesis in the scenarios that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10) and generated from the RGE effects, respectively. For the former scenario, the left panel of Figure 8 has shown the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $2\mu/\Gamma_{\rm tot}$. We see that the results in these five cases are very similar to one another, and the CP asymmetries can take their maximally allowed values around $2\mu/\Gamma_{\rm tot}\sim 1$. This will be helpful for us to reproduce the observed value of $\eta_{\rm B}$. The right panel of Figure 8 has shown the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$. The results for the five cases are very similar to one another, and the ratios of $\mu$ to $M_{0}$ are around $10^{-9}$ or $10^{-18}$ at $M_{0}=10^{3}$ GeV. For the latter scenario, the left panel of Figure 9 has shown the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $M_{0}$, while the right panel of Figure 9 has shown the generated $\eta_{\rm B}$ as functions of $M_{0}$. These results have also been obtained by taking $\Lambda_{\rm FS}=10^{16}$ GeV. The results show that only for Case E and $M_{0}\sim 10^{12}$ GeV can the observed value of $\eta_{\rm B}$ be successfully reproduced, while the the magnitudes of $\eta_{\rm B}$ are always larger than its observed value in the other four cases. It is interesting to recall that Case E holds in the NO case which is slightly preferred by the current experimental results compared to the IO case [2, 3].
Figure 8: For the modular ${\rm S}_{4}$ symmetry model given in section 4, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $2\mu/\Gamma_{\rm tot}$ (left panel), and the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ (right panel).
Figure 9: For the modular ${\rm S}_{4}$ symmetry model given in section 4, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $M_{0}$ (left panel), and the generated $\eta_{\rm B}$ as functions of $M_{0}$ (right panel).
5 Study for modular ${\rm A}_{5}$ symmetry model
In Ref. [24], the authors have presented a comprehensive analysis of neutrino masses and mixing in theories with modular ${\rm A}_{5}$ symmetry. ${\rm A}_{5}$ is the symmetry group of even permutations of 5 objects, and it has one singlet representation ${\bf 1}$, two three-dimensional representations ${\bf 3}$ and ${\bf 3}^{\prime}$, one four-dimensional representation ${\bf 4}$ and one five-dimensional representation ${\bf 5}$. In that paper, the authors have constructed modular-invariant models in which all leptons are collected into ${\bf 3}$ or ${\bf 3}^{\prime}$ multiplets of ${\rm A}_{5}$, containing the three generations of each type of field. The neutrino sector has been taken to be minimal and it only depends on the modulus $\tau$ and an overall scale. But modular forms do not appear in the charged-lepton sector, which instead contains two extra flavons. In the case that the three right-handed neutrinos have a weight $k_{N^{c}}=0$ and constitute the ${\bf 3}$ representation of ${\rm A}_{5}$, the right-handed neutrino mass matrix will take a same form as in Eq. (8). When the three left-handed lepton doublets have a weight $k_{L}=-2$ and constitute ${\bf 3}^{\prime}$ representation of ${\rm A}_{5}$, the Dirac neutrino mass matrix will take a form as
$M_{\rm D}=y_{0}v_{u}\left(\begin{array}[]{ccc}\sqrt{3}Y_{1}&Y_{4}&Y_{3}\\ Y_{5}&-\sqrt{2}Y_{3}&-\sqrt{2}Y_{2}\\ Y_{2}&-\sqrt{2}Y_{5}&-\sqrt{2}Y_{4}\end{array}\right)\;.$ (40)
On the other hand, for the minimal (in terms of weights) possibility, the charged-lepton mass matrix takes a form as
$M_{l}=v_{d}\left(\begin{array}[]{ccc}\alpha+4\gamma(1-\varphi_{2}\varphi_{3})&(\beta+6\gamma)\varphi_{3}&(-\beta+6\gamma)\varphi_{2}\\ (-\beta+6\gamma)\varphi_{3}&6\gamma\varphi_{3}^{2}&\alpha+\beta-2\gamma(1-\varphi_{2}\varphi_{3})\\ (\beta+6\gamma)\varphi_{2}&\alpha-\beta-2\gamma(1-\varphi_{2}\varphi_{3})&6\gamma\varphi_{2}^{2}\end{array}\right)~{}~{}~{}.$ (41)
In the above two equations, $y_{0}$, $\alpha$, $\beta$ and $\gamma$ are simply dimensionless parameters, $\varphi_{2}$ and $\varphi_{3}$ are the VEVs of relevant flavon fields and will be treated as free parameters, while $Y_{i}$ are modular forms of weight 2 and level 5 whose explicit expressions have been formulated in Ref. [25].
Similar to the models in sections 3 and 4, in the present model the charged-lepton masses are essentially controlled by $\alpha v_{d}$, $\beta/\alpha$ and $\gamma/\alpha$, while the neutrino masses and mixing angles are mainly governed by $y^{2}_{0}v^{2}_{u}/M_{0}$, $\tau$, $\varphi_{2}$ and $\varphi_{3}$. The number of parameters is further reduced by enforcing the CP conservation which restricts all the Yukawa couplings together with $\varphi_{2}$ and $\varphi_{3}$ to be real (with the VEV of $\tau$ serving as the only source of CP violation). The numerical calculation shows that this model can be consistent with the neutrino oscillation experimental only in the NO case and it further predicts $m_{1}=0$. The resulting predictions for the best-fit values of the input parameters and observables are collected in Table 5.
Re($\tau$) Im$(\tau)$ $\varphi_{2}$ $\varphi_{3}$ $\alpha\cos\beta\cdot 10^{3}$ $\beta/\alpha$ $\gamma/\alpha$ $y^{2}_{0}\sin^{2}\beta$/$M_{0}$ (eV-1) --0.3615 0.2412 0.04759 0.3731 3.368 1.310 --0.2413 0.0001639
$m_{1}\cdot 10^{2}~{}{\rm eV}^{-1}$ $m_{2}\cdot 10^{2}~{}{\rm eV}^{-1}$ $m_{3}\cdot 10^{2}~{}{\rm eV}^{-1}$ $\sin^{2}\theta_{12}$ $\sin^{2}\theta_{13}$ $\sin^{2}\theta_{23}$ $\delta/\pi$ $(\alpha_{21}$-$\alpha_{31})/\pi$ $\chi^{2}_{\textrm{min}}$ 0 0.860 5.02 0.292 0.0228 0.449 1.63 1.68 11.1
Table 5: For the model given in section 5, in the phenomenologically viable NO case, the predictions for the best-fit values of the input parameters and observables. These results are taken from Tables 8 in Ref. [24].
For this model, in a similar way as in sections 3 and 4, we study the consequences of it for tri-resonant leptogenesis in the scenarios that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10) and generated from the RGE effects, respectively. For the former scenario, the left panel of Figure 10 has shown the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$. The results show that the ratio of $\mu$ to $M_{0}$ is around $10^{-8}$ or $10^{-10}$ at $M_{0}=10^{3}$ GeV. For the latter scenario, the right panel of Figure 10 has shown the generated $\eta_{\rm B}$ as functions of $M_{0}$. These results have also been obtained by taking $\Lambda_{\rm FS}=10^{16}$ GeV. The results show that the observed value of $\eta_{\rm B}$ can not be successfully reproduced.
Figure 10: For the modular ${\rm A}_{5}$ symmetry model given in section 5, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ (left panel). In the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the generated $\eta_{\rm B}$ as functions of $M_{0}$ (right panel).
6 Summary
As we know, due to the peculiarity of the observed neutrino mixing pattern, there have been intensive model-building studies on neutrino-related physics with the employment of certain discrete non-Abelian symmetries. However, a shortcoming of the conventional flavor-symmetry models is that many flavon fields need to be introduced in order to achieve a desired breaking pattern of the flavor symmetry, and the scalar potential of these flavon fields needs a complicated symmetric shaping. In recent years, the idea of modular invariance which provides substantial simplifications to these problems has been proposed and become increasingly popular.
However, for some typical modular symmetry neutrino models, the right-handed neutrino mass matrix takes a form as shown in Eq. (8) which gives three degenerate right-handed neutrino masses and consequently prohibits the leptogenesis to work. But this kind of models can serve as a unique basis for realizations of the tri-resonant leptogenesis scenario: if the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent, then the tri-resonant leptogenesis scenario will be naturally realized. In this paper, we have considered two minimal ways to generate the desired small mass splittings among the three right-handed neutrinos: one way is to modify $M_{\rm R}$ to a form as shown in Eq. (10), while another way is to consider the renormalization-group corrections for the right-handed neutrino masses.
To be specific, we have considered modular ${\rm A}_{4}$, ${\rm S}_{4}$ and ${\rm A}_{5}$ symmetry models that have a right-handed neutrino mass matrix as shown in Eq. (8). For these models, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq. (10), the observed value of $\eta_{\rm B}$ can always be successfully reproduced for appropriate values of $\mu$. These results can be easily understood from the fact that given that $\mu$ is a free parameter, the resonance conditions which enhance the CP asymmetries to their maximally allowed values can always be fulfilled (this is of course helpful for us to reproduce the observed value of $\eta_{\rm B}$).
However, for the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the results will be highly restrictive. Having fixed the flavor-symmetry scale to be at the GUT scale, we reach the following conclusions for the considered representative models:
  • For the modular ${\rm A}_{4}$ symmetry model given in section 3, the observed value of $\eta_{B}$ can be successfully reproduced for $M_{0}\simeq 6\times 10^{12}$ GeV ($4\times 10^{12}$ GeV or $6\times 10^{12}$ GeV) in the NO (IO) case.
  • For the modular ${\rm S}_{4}$ symmetry model given in section 4, among the five cases, only for Case E (which holds in the NO case) and $M_{0}\sim 10^{12}$ GeV can the observed value of $\eta_{\rm B}$ be successfully reproduced.
  • For the modular ${\rm A}_{5}$ symmetry model given in section 5, the observed value of $\eta_{\rm B}$ can not be successfully reproduced.
Acknowledgments
This work is supported in part by the National Natural Science Foundation of China under grant Nos. 11605081, 12142507 and 12147214, the Natural Science Foundation of the Liaoning Scientific Committee under grant NO. 2022-MS-314, and the Ph. D. Research Start-up Fund of Anyang Institute of Technology (no. BSJ2021048).

### v3

@@
-Tri-resonant leptogenesis from modular symmetry neutrino models
+Renormalization group running triggered tri-resonant leptogenesis within neutrino flavor-symmetry models
 Abstract
-In this paper, we have studied the consequences of some representative modular symmetry neutrino models for tri-resonant leptogenesis (which is realized by having three nearly degenerate right-handed neutrinos).
-To be specific, we have considered modular ${\rm A}_{4}$, ${\rm S}_{4}$ and ${\rm A}_{5}$ symmetry models that have a right-handed neutrino mass matrix $M_{\rm R}$ as shown in Eq.
-(8) which gives three degenerate right-handed neutrino masses and consequently prohibits the leptogenesis mechanism to work.
-For these models, we have considered two minimal ways to generate the desired small mass splittings among the three right-handed neutrinos: one scenario is to modify $M_{\rm R}$ to a form as shown in Eq.
-(10), while another scenario is to consider the renormalization-group corrections for the right-handed neutrino masses.
-In the former scenario, for the considered models, the observed value of $\eta_{\rm B}$ can always be successfully reproduced for appropriate values of $\mu$.
-But in the latter scenario not all the considered models can successfully reproduce the observed value of $\eta_{\rm B}$.
+In this paper, we have studied the consequences of some representative neutrino flavor-symmetry models for tri-resonant leptogenesis (which is realized by having three nearly degenerate right-handed neutrinos).
+To be specific, we have considered a neutrino flavor-symmetry model realizing the TM1 mixing and modular ${\rm A}_{4}$, ${\rm S}_{4}$ and ${\rm A}_{5}$ symmetry models that have a right-handed neutrino mass matrix $M_{\rm R}$ as shown in Eq.
+(15) which gives three exactly degenerate right-handed neutrino masses and consequently prohibits the leptogenesis mechanism to work successfully.
+For these models, we study the scenario that the desired right-handed neutrino mass splittings for leptogenesis to work are generated from the renormalization-group running effects.
+In such a scenario, we explore the parameter space that allows for the reproduction of the observed baryon-antibaryon asymmetry of the Universe.
 1 Introduction
-As we know, the discovery of neutrino oscillations reveals that neutrinos are massive and their flavor eigenstates $\nu_{\alpha}$ (for $\alpha=e,\mu,\tau$) are certain superpositions of their mass eigenstates $\nu_{i}$ (for $i=1,2,3$) possessing definite masses $m_{i}$: $\nu_{\alpha}=\sum_{i}U_{\alpha i}\nu_{i}$ with $U_{\alpha i}$ being the $\alpha i$-th element of the neutrino mixing matrix $U$ [1].
+As we know, the discovery of neutrino oscillations reveals that neutrinos are massive and their flavor eigenstates $\nu_{\alpha}$ (for $\alpha=e,\mu,\tau$) are certain superpositions of their mass eigenstates $\nu_{i}$ (for $i=1,2,3$) possessing definite masses $m_{i}$: $\nu_{\alpha}=\sum_{i}U_{\alpha i}\nu_{i}$ with $U_{\alpha i}$ being the $\alpha i$-element of the neutrino mixing matrix $U$ [1].
 In the standard form, $U$ is expressed in terms of three mixing angles $\theta_{ij}$ (for $ij=12,13,23$), one Dirac CP phase $\delta$ and two Majorana CP phases $\alpha_{21}$ and $\alpha_{31}$ as
@@
 Nowadays, the neutrino mixing angles and neutrino mass squared differences $\Delta m^{2}_{ij}\equiv m^{2}_{i}-m^{2}_{j}$ have been measured to good degrees of accuracy, and there is also a preliminary result for $\delta$ (but with a large uncertainty).
-Several research groups have performed global analyses of the existing neutrino oscillation data to extract the values of these parameters [2, 3].
-For reference, the results in Ref.
-[2] are reproduced in Table 1 here.
-Note that the sign of $\Delta m^{2}_{31}$ remains undetermined, thereby allowing for two possible neutrino mass orderings: the normal ordering (NO) $m_{1}<m_{2}<m_{3}$ and inverted ordering (IO) $m_{3}<m_{1}<m_{2}$.
-However, neutrino oscillations are completely insensitive to the absolute neutrino mass scale and the Majorana CP phases.
+Several research groups have performed global analyses of the existing neutrino oscillation data to extract the values of these parameters [2].
+As is known, the sign of $\Delta m^{2}_{31}$ remains undetermined, thereby allowing for two possible neutrino mass orderings: the normal ordering (NO) $m_{1}<m_{2}<m_{3}$ and inverted ordering (IO) $m_{3}<m_{1}<m_{2}$.
+On the other hand, neutrino oscillations are completely insensitive to the absolute neutrino mass scale and the Majorana CP phases.
 Their values can only be inferred from certain non-oscillatory experiments such as the neutrinoless double beta decay experiments [3].
 But so far there has not been any lower bound on the lightest neutrino mass, nor any constraint on the Majorana CP phases.
-For the neutrino mixing angles, Table 1 shows that $\theta_{12}$ and $\theta_{23}$ are close to some special values (i.e., $\sin^{2}\theta_{12}\sim 1/3$ and $\sin^{2}\theta_{23}\sim 1/2$).
-But $\theta_{13}$ is relatively small.
-In fact, before its value was measured, $\theta_{13}$ had been widely expected to be vanishingly small.
+For the neutrino mixing angles, it is known that $\theta_{12}$ and $\theta_{23}$ are close to some special values (i.e., $\sin^{2}\theta_{12}\sim 1/3$ and $\sin^{2}\theta_{23}\sim 1/2$), while $\theta_{13}$ is relatively small.
+In fact, before the true value of $\theta_{13}$ was experimentally pinned down, it had been widely believed to be vanishingly small.
 For the ideal case of $\sin^{2}\theta_{12}=1/3$, $\sin^{2}\theta_{23}=1/2$ and $\theta_{13}=0$, the neutrino mixing matrix takes a very simple form as
@@
 This particular mixing has inspired intensive model-building studies with the employment of certain discrete non-Abelian symmetries (e.g., ${\rm A}_{4}$ and ${\rm S}_{4}$) [5].
-Noteworthy, a shortcoming of the conventional flavor-symmetry models is that many flavon fields need to be introduced in order to achieve a desired breaking pattern of the flavor symmetry, and the scalar potential of these flavon fields needs a complicated symmetric shaping.
+However, the observation of a relatively large $\theta_{13}$ (compared to 0) motivates us to make corrections for the TBM mixing.
+In this connection, an economical and natural choice is to retain its first or second column while modifying the other two columns within the unitary constraints, thus yielding the first or second trimaximal (TM1 or TM2) mixing [6]
+$U_{\rm TM1}=\displaystyle\frac{1}{\sqrt{6}}\left(\begin{array}[]{ccc}2&\cdot&\cdot\cr 1&\cdot&\cdot\cr 1&\cdot&\cdot\cr\end{array}\right)\;,\hskip 28.45274ptU_{\rm TM2}=\displaystyle\frac{1}{\sqrt{3}}\left(\begin{array}[]{ccc}\cdot&1&\cdot\cr\cdot&-1&\cdot\cr\cdot&-1&\cdot\cr\end{array}\right)\;,$ (12)
+where the dot signs denote the unspecified elements.
+Due to their simplicity and interesting consequences, the TM1 and TM2 mixings have attracted a lot of attention and many neutrino flavor-symmetry models have been formulated to realize them [5].
+An obvious shortcoming of the conventional flavor-symmetry models is that many flavon fields need to be introduced in order to achieve a desired breaking pattern of the flavor symmetry, and the scalar potential of these flavon fields needs a complicated symmetric shaping.
 Moreover, higher-dimensional operators with more flavon insertions and unknown coefficients are often present, affecting the predictive power of the models.
 In recent years, the idea of modular invariance which provides substantial simplifications to these problems has been proposed and become increasingly popular [7, 8].
-In the modular-invariance approach, the flavour symmetry is broken by the VEV of a single scalar field—the so-called modulus, while the flavon fields are not necessary.
+In the modular-invariance approach, the flavor symmetry is broken by the vacuum expectation value (VEV) of a single scalar field—the so-called modulus, while the flavon fields are not necessary.
 And the Yukawa couplings and fermion mass terms are certain special functions of the modulus called modular forms that have particular properties under the modular group transformations.
@@
 As for the neutrino masses, one of the most popular and natural ways of generating them is the type-I seesaw model [9] in which three heavy right-handed neutrinos $N_{I}$ ($I=1,2,3$) are introduced into the Standard Model (SM).
-First of all, $N_{I}$ can constitute the Yukawa coupling operators together with the lepton doublets $L_{\alpha}$ and the Higgs doublet $H$: $Y_{I\alpha}\overline{N_{I}}HL_{\alpha}$ with $Y_{I\alpha}$ being the $I\alpha$-th element of the Yukawa coupling matrix $Y$.
-These operators will generate the Dirac neutrino mass terms $(M_{\rm D})_{I\alpha}=Y_{I\alpha}v$ [here $(M_{\rm D})_{I\alpha}$ is the $I\alpha$-th element of the Dirac neutrino mass matrix $M_{\rm D}$] after the neutral component of $H$ acquires a nonzero vacuum expectation value (VEV) $v=174$ GeV.
-In addition, $N_{I}$ themselves can also have the Majorana mass terms $\overline{N_{I}}(M_{\rm R})_{IJ}N^{c}_{J}$ [here $(M_{\rm R})_{IJ}$ are the $IJ$-th element of the right-handed neutrino mass matrix $M_{\rm R}$].
+First of all, $N_{I}$ can constitute the Yukawa coupling operators together with the lepton doublets $L_{\alpha}$ and the Higgs doublet $H$: $Y_{I\alpha}\overline{N_{I}}HL_{\alpha}$ with $Y_{I\alpha}$ being the $I\alpha$-element of the Yukawa coupling matrix $Y$.
+These operators will generate the Dirac neutrino mass terms $(M_{\rm D})_{I\alpha}=Y_{I\alpha}v$ (here $(M_{\rm D})_{I\alpha}$ is the $I\alpha$-element of the Dirac neutrino mass matrix $M_{\rm D}$) after the neutral component of $H$ acquires a nonzero VEV $v=174$ GeV.
+In addition, $N_{I}$ themselves can also have the Majorana mass terms $\overline{N_{I}}(M_{\rm R})_{IJ}N^{c}_{J}$ (here $(M_{\rm R})_{IJ}$ is the $IJ$-element of the right-handed neutrino mass matrix $M_{\rm R}$).
 Then, under the seesaw condition $M_{\rm R}\gg M_{\rm D}$, integrating the right-handed neutrinos out yields an effective Majorana mass matrix for three light neutrinos as
 $M_{\nu}=-M^{T}_{\rm D}M^{-1}_{\rm R}M_{\rm D}\;.$ (13)
-Here we would like to point out that throughout the paper the fermion mass matrices are written in the right-left convention which is more conveniently used in the supersymmetric setup as applicable to our study.
-Normal Ordering Inverted Ordering
-bfp ±$\pm 1\sigma$ $3\sigma$ range bfp ±$\pm 1\sigma$ $3\sigma$ range
-$\sin^{2}\theta_{12}$ $0.303_{-0.011}^{+0.012}$ $0.270\to 0.341$ $0.303_{-0.011}^{+0.012}$ $0.270\to 0.341$
-$\sin^{2}\theta_{23}$ $0.572_{-0.023}^{+0.018}$ $0.406\to 0.620$ $0.578_{-0.021}^{+0.016}$ $0.412\to 0.623$
-$\sin^{2}\theta_{13}$ $0.02203_{-0.00059}^{+0.00056}$ $0.02029\to 0.02391$ $0.02219_{-0.00057}^{+0.00060}$ $0.02047\to 0.02396$
-$\delta/\pi$ $1.09_{-0.14}^{+0.23}$ $0.96\to 1.33$ $1.59_{-0.18}^{+0.15}$ $1.41\to 1.74$
-$\Delta m^{2}_{21}/(10^{-5}~{}{\rm eV}^{2})$ $7.41_{-0.20}^{+0.21}$ $6.82\to 8.03$ $7.41_{-0.20}^{+0.21}$ $6.82\to 8.03$
-$\Delta m^{2}_{3\ell}/(10^{-3}~{}{\rm eV}^{2})$ $2.511_{-0.027}^{+0.028}$ $2.428\to 2.597$ $-2.498_{-0.025}^{+0.032}$ $-2.581\to-2.408$
-[…]
-Remarkably, the seesaw model also provides an attractive explanation (which is known as the leptogenesis mechanism [10, 11]) for the baryon-antibaryon asymmetry of the Universe [12]
-$Y_{\rm B}\equiv\frac{n_{\rm B}-n_{\rm\bar{B}}}{s}\simeq(8.69\pm 0.04)\times 10^{-11}\;,$ (7)
-where $n_{\rm B}$ ($n_{\rm\bar{B}}$) denotes the baryon (antibaryon) number density and $s$ the entropy density.
-The leptogenesis mechanism works in a way as follows: a lepton-antilepton asymmetry $Y_{\rm L}\equiv(n_{\rm L}-n_{\rm\bar{L}})/s$ is firstly generated from the out-of-equilibrium and CP-violating decays of the right-handed neutrinos and then partly converted into the baryon-antibaryon asymmetry via the sphaleron processes: $Y_{\rm B}\simeq-cY_{\rm L}$ with $c=28/79$ or $8/23$ in the SM or MSSM (Minimal Supersymmetric Standard Model) framework.
-As is known, in the scenario that the right-handed neutrino masses are hierarchial there exists a lower bound about $10^{9}$ GeV for the right-handed neutrino mass scale in order to successfully reproduce the observed value of $Y_{\rm B}$ [13].
-Obviously, such a high right-handed neutrino mass scale has no chance to be directly accessed by forseeable collider experiments.
-Fortunately, in the resonant leptogenesis scenario which is realized for nearly degenerate right-handed neutrinos [14], a successful leptogenesis can be realized even with TeV-scale right-handed neutrinos which have the potential to be directly accessed by running or upcoming experiments [15].
-However, for some typical modular symmetry neutrino models (see the models provided in sections 3-5), the right-handed neutrino mass matrix takes a form as
-$M_{\rm R}=\begin{pmatrix}M_{0}&0&0\\ 0&0&M_{0}\\ 0&M_{0}&0\end{pmatrix}\;.$ (8)
-It can be diagonalized by the following unitary matrix
-$U_{\rm R}=\begin{pmatrix}0&1&0\\ \frac{\rm i}{\sqrt{2}}&0&\frac{1}{\sqrt{2}}\\ -\frac{\rm i}{\sqrt{2}}&0&\frac{1}{\sqrt{2}}\end{pmatrix}\;,$ (9)
-to give three degenerate right-handed neutrino masses (i.e., $M_{0}$).
-In this case, unfortunately, the leptogenesis mechanism is prohibited to work.
-But just as a coin has two sides, this kind of models can serve as a unique basis for realizations of the tri-resonant leptogenesis scenario: if the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent, then the tri-resonant leptogenesis scenario will be naturally realized For a realization of resonant leptogenesis in the minimal seesaw model with a modular symmetry, see Ref.
-[16].
-.
-In this paper, we consider two minimal ways to generate the desired small mass splittings among the three right-handed neutrinos: one way is to modify $M_{\rm R}$ to a form as
-$M_{\rm R}=\begin{pmatrix}M_{0}&0&0\\ 0&\mu&M_{0}\\ 0&M_{0}&\mu\end{pmatrix}\;.$ (10)
-with $\mu\ll M_{0}$.
-Such a modification of $M_{\rm R}$ has the merit that it only introduces one additional parameter and it still can be diagonalized by the unitary matrix $U_{\rm R}$ in Eq.
-(9) (i.e., it does not alter the original neutrino mixing pattern).
-But the three right-handed neutrino masses are modified to $M_{0}-\mu$, $M_{0}$ and $M_{0}+\mu$, respectively.
-Another way is to consider the renormalization-group corrections for the right-handed neutrino masses [17]: if the energy scale $\Lambda_{\rm FS}$ of the flavor-symmetry physics that shapes the special textures of neutrino mass matrices is much higher than the right-handed neutrino mass scale $M_{0}$ where leptogenesis takes place (such a scenario is very likely given that the idea of modular flavor symmetry is just inspired from the string theory [7]), then the renormalization group evolution (RGE) effect between the scales of $\Lambda_{\rm FS}$ and $M_{0}$ may induce some mass splittings among the three right-handed neutrinos (see section 2.2).
+Here we would like to emphasize that throughout the paper the fermion mass matrices are written in the right-left convention which is more conveniently used in the supersymmetric setup as applicable to our study.
+Remarkably, the seesaw model also provides an attractive explanation (known as the leptogenesis mechanism [10, 11]) for the baryon-antibaryon asymmetry of the Universe [12]
+$\eta_{\rm B}\equiv\frac{n_{\rm B}-n_{\rm\bar{B}}}{n_{\gamma}}\simeq(6.11\pm 0.04)\times 10^{-10}\;,$ (14)
+where $n_{\rm B}$ ($n_{\rm\bar{B}}$) denotes the baryon (antibaryon) number density and $n_{\gamma}$ the photon number density.
+The leptogenesis mechanism works in a way as follows: a lepton-antilepton asymmetry $\eta_{\rm L}\equiv(n_{\rm L}-n_{\rm\bar{L}})/n_{\gamma}$ is firstly generated from the out-of-equilibrium and CP-violating decays of the right-handed neutrinos and then partly converted into the baryon-antibaryon asymmetry via the sphaleron processes: $\eta_{\rm B}\simeq-c\eta_{\rm L}$ with $c=28/79$ or $8/23$ in the SM or MSSM (Minimal Supersymmetric Standard Model) framework.
+Unfortunately, for some typical neutrino flavor-symmetry models (see those given in sections 3-6), the right-handed neutrino mass matrix takes a form as
+$M_{\rm R}=M_{0}\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix}\;$, (15)
+which gives three exactly degenerate right-handed neutrino masses (i.e., $M_{0}$).
+In this case, the leptogenesis mechanism is prohibited to work successfully (to be specific, there would no CP asymmetries for the decays of right-handed neutrinos).
+But just as a coin has two sides, this kind of models can serve as a unique basis for realizations of the tri-resonant leptogenesis scenario [13]: if the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent, then the tri-resonant leptogenesis scenario will be naturally realized.
+In this paper, within this kind of neutrino flavor-symmetry models, we study the scenario that the desired right-handed neutrino mass splittings for leptogenesis to work are generated from the renormalization-group equation (RGE) effects [14]: if the energy scale $\Lambda_{\rm FS}$ of the flavor-symmetry physics that shapes the special textures of neutrino mass matrices is much higher than the right-handed neutrino mass scale $M_{0}$ (throughout this paper, $M_{0}$ will be used to denote the common masses of three nearly degenerate right-handed neutrinos when their splitting is not important) where leptogenesis takes place, then the RGE effect between the scales of $\Lambda_{\rm FS}$ and $M_{0}$ may induce some mass splittings among the three right-handed neutrinos (see section 2.2).
 Such a correction of $M_{\rm R}$ has the following two merits: 1) this effect is spontaneous, provided that there is a considerable gap between the flavor-symmetry scale and the right-handed neutrino mass scale; 2) this effect is minimal, in the sense that it does not need to introduce additional parameters.
 The remaining parts of this paper are organized as follows.
-In section 2 we will first give some basic formulas about resonant leptogenesis and RGE effects that will be used in our study.
-Then, we perform the study for three representative modular symmetry models which contain three degenerate right-handed neutrinos.
-We first consider the modular ${\rm A}_{4}$ symmetry model in section 3, then the modular ${\rm S}_{4}$ symmetry model in section 4, and finally the modular ${\rm A}_{5}$ symmetry model in section 5.
+In section 2 we will first give some basic formulas about leptogenesis and RGE effects that will be used in our study.
+Then, we perform the study for four representative neutrino flavor-symmetry models which contain three degenerate right-handed neutrinos For a similar study about the $\Delta(3n^{2})$ and $\Delta(6n^{2})$ neutrino flavor-symmetry models, see Ref.
+[15]..
+We successively study a neutrino flavor-symmetry model realizing the TM1 mixing in section 3, modular ${\rm A}_{4}$ symmetry model in section 4, modular ${\rm S}_{4}$ symmetry model in section 5, and modular ${\rm A}_{5}$ symmetry model in section 6.
 In section 7, we summarize our main results.
-2 Resonant leptogenesis and RGE effects
-2.1 Some basics about resonant leptogenesis
-As is known, if the temperature $T$ of the Universe at which leptogenesis takes place (approximately the right-handed neutrino mass scale) is below $10^{9}$ GeV, all the three lepton flavors have become distinguishable from one another and the lepton asymmetries stored in them should be tracked separately [18].
-Such a three-flavor regime applies to the scenario studied in this paper (i.e., low-scale tri-resonant leptogenesis).
-In this case, the final lepton asymmetry can be approximately expressed as
-$Y_{\rm L}\simeq r(\varepsilon_{e}\kappa_{e}+\varepsilon_{\mu}\kappa_{\mu}+\varepsilon_{\tau}\kappa_{\tau})\;.$ (11)
-Here the factor $r\simeq 4\times 10^{-3}$ measures the ratio of the equilibrium number density of right-handed neutrinos to the entropy density.
-And $\varepsilon_{e}$ is a sum over three right-handed neutrinos (i.e., $\varepsilon_{e}=\varepsilon_{1e}+\varepsilon_{2e}+\varepsilon_{3e}$) of the following CP asymmetries:
-$\varepsilon_{I\alpha}\equiv\frac{\Gamma\left(N_{I}\to L_{\alpha}+H\right)-\Gamma\left(N_{I}\to\overline{L}_{\alpha}+\overline{H}\right)}{\sum_{\alpha}\Gamma\left(N_{I}\to L_{\alpha}+H\right)+\Gamma\left(N_{I}\to\overline{L}_{\alpha}+\overline{H}\right)}\;,$ (12)
-whose explicit expressions will be given below.
-Finally, $\kappa_{\alpha}$ are the efficiency factors (which are smaller than 1) due to the washout effects.
-Their concrete values can be calculated by numerically solving relevant Boltzmann equations, whose explicit expressions will also be given below.
-To be explicit, in the basis where both the right-handed neutrino matrix and charged lepton mass matrix are diagonal, the CP asymmetries $\varepsilon_{I\alpha}$ defined in Eq.
-(12) can be expressed as [14]
-$\varepsilon_{I\alpha}=\frac{\left|(\bar{Y}_{+})_{I\alpha}\right|^{2}-\left|(\bar{Y}_{-})_{I\alpha}\right|^{2}}{\left(\bar{Y}_{+}\bar{Y}^{\dagger}_{+}\right)_{II}+\left(\bar{Y}_{-}\bar{Y}^{\dagger}_{-}\right)_{II}}\;.$ (13)
-Here $(\bar{Y}_{+})_{I\alpha}$ are the effective neutrino Yukawa couplings in the case of there existing three nearly degenerate right-handed neutrinos. They can be expressed as
-$\left(\bar{Y}_{+}\right)_{I\alpha}=Y_{I\alpha}-{\rm i}V_{I\alpha}+{\rm i}\sum_{J,K=1}^{3}\left|\epsilon_{IJK}\right|Y_{J\alpha}$
-×$\times\frac{M_{I}\left(\mathcal{M}_{IIJ}+\mathcal{M}_{JJI}\right)+{\rm i}R_{IK}\left[\mathcal{M}_{IKJ}\left(\mathcal{M}_{IIK}+\mathcal{M}_{KKI}\right)+\mathcal{M}_{JJK}\left(\mathcal{M}_{IKI}+\mathcal{M}_{KIK}\right)\right]}{M^{2}_{I}-M^{2}_{J}-2{\rm i}M^{2}_{I}A_{JJ}-2{\rm i}{\rm Im}(R_{IK})\left[M^{2}_{I}\left|A_{JK}\right|^{2}+M_{J}M_{K}{\rm Re}(A^{2}_{JK})\right]}\;,$ (14)
+2 Some formulas about leptogenesis and RGE effects
+2.1 Some formulas about leptogenesis
+In the resonant leptogenesis scenario, the CP asymmetries for the decay processes of right-handed neutrinos will be greatly enhanced due to nearly degenerate masses of the right-handed neutrinos.
+The resonance effects can be encoded in the effective neutrino Yukawa couplings as [13]
+$\hskip-56.9055pt\left(\bar{Y}_{+}\right)_{I\alpha}=Y_{I\alpha}-{\rm i}V_{I\alpha}+{\rm i}\sum_{J,K=1}^{3}\left|\epsilon_{IJK}\right|Y_{J\alpha}$
+$\hskip-5.69046pt\times\frac{M_{I}\left(\mathcal{M}_{IIJ}+\mathcal{M}_{JJI}\right)+{\rm i}R_{IK}\left[\mathcal{M}_{IKJ}\left(\mathcal{M}_{IIK}+\mathcal{M}_{KKI}\right)+\mathcal{M}_{JJK}\left(\mathcal{M}_{IKI}+\mathcal{M}_{KIK}\right)\right]}{M^{2}_{I}-M^{2}_{J}-2{\rm i}M^{2}_{I}A_{JJ}-2{\rm i}{\rm Im}(R_{IK})\left[M^{2}_{I}\left|A_{JK}\right|^{2}+M_{J}M_{K}{\rm Re}(A^{2}_{JK})\right]}\;,$ (16)
 where $\epsilon_{IJK}$ is the anti-symmetric Levi-Civita symbol, $\mathcal{M}_{IJK}\equiv M_{I}A_{JK}$ and
-[…]
+$A_{IJ}=\frac{\left(YY^{\dagger}\right)^{*}_{IJ}}{16\pi}\;,\hskip 28.45274ptV_{I\alpha}=\sum_{\beta=e,\mu,\tau}\sum_{K\neq I}\frac{Y^{*}_{I\beta}Y_{K\beta}Y_{K\alpha}}{16\pi}f\left(\frac{M^{2}_{K}}{M^{2}_{I}}\right)\;,$
 $R_{IJ}=\frac{M^{2}_{I}}{M^{2}_{I}-M^{2}_{J}-2{\rm i}M^{2}_{I}A_{JJ}}\;,$ (17)
@@
 And the corresponding CP-conjugate effective Yukawa couplings $\left(\bar{Y}_{-}\right)_{I\alpha}$ can be obtained by simply making the replacement $Y_{I\alpha}\to Y^{*}_{I\alpha}$ in the above expressions.
-Furthermore, in terms of these effective Yukawa couplings, the decay widths of right-handed neutrinos can be expressed as
-$\Gamma\left(N_{I}\to L_{\alpha}+H\right)=\frac{M_{I}}{16\pi}\left|\left(\bar{Y}_{+}\right)_{I\alpha}\right|^{2}\;,\hskip 28.45274pt\Gamma\left(N_{I}\to\overline{L}_{\alpha}+\overline{H}\right)=\frac{M_{I}}{16\pi}\left|\left(\bar{Y}_{-}\right)_{I\alpha}\right|^{2}\;,$ (16)
-and their branching ratios are given by
-$B_{I\alpha}=\frac{\left|(\bar{Y}_{+})_{I\alpha}\right|^{2}+\left|(\bar{Y}_{-})_{I\alpha}\right|^{2}}{\left(\bar{Y}_{+}\bar{Y}^{\dagger}_{+}\right)_{II}+\left(\bar{Y}_{-}\bar{Y}^{\dagger}_{-}\right)_{II}}\;.$ (17)
-In order to generate a nonzero lepton asymmetry in the Universe, one needs not only the above CP asymmetries, but also a departure from thermal equilibrium.
-In the latter processes, the evolutions of the number densities of the right-handed neutrinos $n_{N_{I}}$ and flavored lepton asymmetries $n_{L_{\alpha}}$ (which are normalized by the photon number density $n_{\gamma}$ to $\eta_{N_{I}}\equiv n_{N_{I}}/n_{\gamma}$ and $\eta_{L_{\alpha}}\equiv n_{L_{\alpha}}/n_{\gamma}$, respectively) can be tracked by the following Boltzmann equations [14]
-$\frac{{\rm d}\eta_{N_{I}}}{{\rm d}z}$ $=$ $\frac{z}{H(z=1)}\Bigg{[}\left(1-\frac{\eta_{N_{I}}}{\eta_{N_{I}}^{\mathrm{eq}}}\right)\sum_{\alpha=e,\mu,\tau}\left(\Gamma^{D(I\alpha)}+\Gamma_{\text{Yukawa }}^{S(I\alpha)}+\Gamma_{\text{Gauge }}^{S(I\alpha)}\right)$
-$-\frac{2}{3}\sum_{\alpha=e,\mu,\tau}\eta_{L_{\alpha}}\varepsilon_{I\alpha}\left(\hat{\Gamma}^{D(I\alpha)}+\hat{\Gamma}_{\text{Yukawa }}^{S(I\alpha)}+\hat{\Gamma}_{\text{Gauge }}^{S(I\alpha)}\right)\Bigg{]}\;,$
-$\frac{{\rm d}\eta_{L_{\alpha}}}{{\rm d}z}$ $=$ $\frac{z}{H(z=1)}\Bigg{\{}\sum^{3}_{I=1}\varepsilon_{I\alpha}\left(\frac{\eta_{N_{I}}}{\eta_{N_{I}}^{\mathrm{eq}}}-1\right)\sum_{\beta=e,\mu,\tau}\left(\Gamma^{D(I\beta)}+\Gamma_{\text{Yukawa }}^{S(I\beta)}+\Gamma_{\text{Gauge }}^{S(I\beta)}\right)$ (18)
-$-\frac{2}{3}\eta_{L_{\alpha}}\Bigg{[}\sum^{3}_{I=1}B_{I\alpha}\left(\tilde{\Gamma}^{D(I\alpha)}+\tilde{\Gamma}_{\text{Yukawa }}^{S(I\alpha)}+\tilde{\Gamma}_{\text{Gauge }}^{S(I\alpha)}+\Gamma_{\text{Yukawa }}^{W(I\alpha)}+\Gamma_{\text{Gauge }}^{W(I\alpha)}\right)$
-$\hskip 42.67912pt+\sum_{\beta=e,\mu,\tau}\left(\Gamma_{\text{Yukawa }}^{\Delta L=2(\alpha\beta)}+\Gamma_{\text{Yukawa }}^{\Delta L=0(\alpha\beta)}\right)\Bigg{]}$
-$-\frac{2}{3}\sum_{\beta=e,\mu,\tau}\eta_{L_{\beta}}\Bigg{[}\sum^{3}_{I=1}\varepsilon_{I\alpha}\varepsilon_{I\beta}\left(\Gamma_{\text{Yukawa }}^{W(I\beta)}+\Gamma_{\text{Gauge }}^{W(I\beta)}+\Gamma_{\text{Yukawa }}^{\Delta L=2(\alpha\beta)}-\Gamma_{\text{Yukawa }}^{\Delta L=0(\alpha\beta)}\right)\Bigg{]}\Bigg{\}}\;,$
-where $z=M_{1}/T$ has been defined, and
-$H\left(z\right)=\sqrt{\frac{4\pi^{3}g_{*}}{45}}\frac{M^{2}_{1}}{z^{2}M_{\mathrm{Pl}}}\;,$ (19)
-is the Hubble parameter with $M_{\mathrm{Pl}}=1.221\times 10^{19}\mathrm{GeV}$ and $g_{*}\simeq 106.75$ or 228.75 (in the SM and MSSM frameworks, respectively) being the number of relativistic degrees of freedom in the early Universe.
-And the analytical expressions of the $\Gamma$ terms on the right-hand side of Eq.
-(18) can be found in Ref.
+In terms of these effective Yukawa couplings, the decay widths of right-handed neutrinos can be expressed as
+$\Gamma_{I}\equiv\sum_{\alpha}\Gamma\left(N_{I}\to L_{\alpha}+H\right)=\frac{M_{I}}{16\pi}\left|\left(\bar{Y}_{+}\bar{Y}^{\dagger}_{+}\right)_{II}\right|^{2}\;,\quad\bar{\Gamma}_{I}\equiv\sum_{\alpha}\Gamma\left(N_{I}\to\overline{L}_{\alpha}+\overline{H}\right)=\frac{M_{I}}{16\pi}\left|\left(\bar{Y}_{-}\bar{Y}^{\dagger}_{-}\right)_{II}\right|^{2}\;,$ (18)
+Finally, the CP-asymmetry matrix element $\varepsilon^{\left(I\right)}_{\alpha\beta}$ describing the asymmetries generated from the decays of $N_{I}$ can be obtained as
+$\varepsilon^{\left(I\right)}_{\alpha\beta}=\frac{\bar{\mathcal{P}}^{\left(I\right)}_{\alpha\beta}\bar{\Gamma}_{I}-\mathcal{P}^{\left(I\right)}_{\alpha\beta}\Gamma_{I}}{\Gamma_{I}+\bar{\Gamma}_{I}}\;.$ (19)
+Here the two projectors $\mathcal{P}^{\left(I\right)}_{\alpha\beta}$ and $\bar{\mathcal{P}}^{\left(I\right)}_{\alpha\beta}$ are given by
+$\mathcal{P}^{\left(I\right)}_{\alpha\beta}\equiv C_{\alpha I}C^{*}_{\beta I}\;,\qquad\bar{\mathcal{P}}^{\left(I\right)}_{\alpha\beta}\equiv\bar{C}_{\alpha I}\bar{C}^{*}_{\beta I}\;,$ (20)
+with
+$C_{I\alpha}=\frac{\left(\bar{Y}_{+}\right)_{I\alpha}}{\sqrt{\left(\bar{Y}_{+}\bar{Y}^{\dagger}_{+}\right)_{II}}}\;,\qquad\bar{C}_{I\alpha}=\frac{\left(\bar{Y}_{-}\right)_{I\alpha}}{\sqrt{\left(\bar{Y}_{-}\bar{Y}^{\dagger}_{-}\right)_{II}}}\;.$ (21)
+Neglecting CP-violating corrections, $C_{\alpha I}$ and $\bar{C}_{\alpha I}$ reduce to
+$C^{0}_{I\alpha}=\frac{Y_{I\alpha}}{\sqrt{\left(YY^{\dagger}\right)_{II}}}\;.$ (22)
+[…]
+In order to generate a nonzero lepton asymmetry in the Universe, one needs not only the CP asymmetries, but also a departure from thermal equilibrium.
+In most cases, classical Boltzmann equations are sufficient for the calculation of the final asymmetry [11].
+However, when lepton flavour effects are taken into account [16], different sets of classical Boltzmann equations apply depending whether the asymmetry is generated in the unflavored regime (where none of the charged-lepton Yukawa interactions has entered into equilibrium), in the two-flavour regime (where only the tauonic Yukawa interaction has entered into equilibrium), or in the three-flavour regime (where both the tauonic and muonic Yukawa interactions have entered into equilibrium).
+In particular, classical Boltzmann equations would fail in reproducing the correct result in the transition regimes between any two different flavor regimes.
+As our study aims to explore the full parameter space of the right-handed neutrino masses, we will employ the density matrix formalism [17] which applies to different flavor regimes as well as the transition regimes.
+In this formalism the evolutions of the right-handed neutrino number density $n_{N_{I}}$ and lepton asymmetry number density $n_{\alpha\beta}$ as functions of the temperature (via $z\equiv M_{1}/T$) can be tracked by the following equations [17]
+$\frac{{\rm d}n_{N_{I}}}{{\rm d}z}$ $=$ $-D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)\;,$
+$\frac{{\rm d}n_{ee}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\left\{\varepsilon^{\left(I\right)}_{ee}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-W_{I}\left[\left|C^{0}_{Ie}\right|^{2}n_{ee}+{\rm Re}\left(C^{0}_{Ie}C^{0}_{I\mu}n^{*}_{\mu e}+C^{0}_{Ie}C^{0}_{I\tau}n^{*}_{\tau e}\right)\right]\right\}\;,$
+$\frac{{\rm d}n_{\mu\mu}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\left\{\varepsilon^{\left(I\right)}_{\mu\mu}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-W_{I}\left[\left|C^{0}_{I\mu}\right|^{2}n_{\mu\mu}+{\rm Re}\left(C^{0}_{I\mu}C^{0}_{Ie}n_{\mu e}+C^{0}_{I\mu}C^{0}_{I\tau}n^{*}_{\tau\mu}\right)\right]\right\}\;,$
+$\frac{{\rm d}n_{\tau\tau}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\left\{\varepsilon^{\left(I\right)}_{\tau\tau}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-W_{I}\left[\left|C^{0}_{I\tau}\right|^{2}n_{\tau\tau}+{\rm Re}\left(C^{0}_{I\tau}C^{0}_{Ie}n_{\tau e}+C^{0}_{I\tau}C^{0}_{I\mu}n_{\tau\mu}\right)\right]\right\}\;,$
+$\frac{{\rm d}n_{\tau\mu}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\bigg\{\varepsilon^{\left(I\right)}_{\tau\mu}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-\frac{1}{2}W_{I}\bigg[n_{\tau\mu}\left(\left|C^{0}_{I\tau}\right|^{2}+\left|C^{0}_{I\mu}\right|^{2}\right)+C^{0*}_{I\mu}C^{0}_{I\tau}\left(n_{\tau\tau}+n_{\mu\mu}\right)$
+$\hskip 28.45274pt+C^{0*}_{Ie}C^{0}_{I\tau}n^{*}_{\mu e}+C^{0*}_{I\mu}C^{0}_{Ie}n_{\tau e}\bigg]\bigg\}-\left[\frac{{\rm Im}\left(\Lambda_{\tau}\right)}{Hz}+\frac{{\rm Im}\left(\Lambda_{\mu}\right)}{Hz}\right]n_{\tau\mu}\;,$
+$\frac{{\rm d}n_{\tau e}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\bigg\{\varepsilon^{\left(I\right)}_{\tau e}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-\frac{1}{2}W_{I}\bigg[n_{\tau e}\left(\left|C^{0}_{Ie}\right|^{2}+\left|C^{0}_{I\tau}\right|^{2}\right)+C^{0*}_{Ie}C^{0}_{I\tau}\left(n_{ee}+n_{\tau\tau}\right)$
+$\hskip 28.45274pt+C^{0*}_{I\mu}C^{0}_{I\tau}n_{\mu e}+C^{0*}_{Ie}C^{0}_{I\mu}n_{\tau\mu}\bigg]\bigg\}-\frac{{\rm Im}\left(\Lambda_{\tau}\right)}{Hz}n_{\tau e}\;,$
+$\frac{{\rm d}n_{\mu e}}{{\rm d}z}$ $=$ $\sum_{I=1}^{3}\bigg\{\varepsilon^{\left(I\right)}_{\mu e}D_{I}\left(n_{N_{I}}-n^{\rm eq}_{N_{I}}\right)-\frac{1}{2}W_{I}\bigg[n_{\mu e}\left(\left|C^{0}_{Ie}\right|^{2}+\left|C^{0}_{I\mu}\right|^{2}\right)+C^{0*}_{Ie}C^{0}_{I\mu}\left(n_{ee}+n_{\mu\mu}\right)$ (23)
+$\hskip 28.45274pt+C^{0*}_{Ie}C^{0}_{I\tau}n^{*}_{\tau\mu}+C^{0*}_{I\tau}C^{0}_{I\mu}n_{\tau e}\bigg]\bigg\}-\frac{{\rm Im}\left(\Lambda_{\mu}\right)}{Hz}n_{\mu e}\;,$
+where $C^{0}_{I\alpha}$ have been defined in Eq.
+(22).
+In these equations, $n^{\rm eq}_{N_{I}}$ are the equilibrium number density of three right-handed neutrinos.
+The decay (washout) parameters $D_{I}\left(W_{I}\right)$ are given by
+$D_{I}\left(z\right)=K_{I}x_{I}z\frac{\mathcal{K}_{1}\left(z_{I}\right)}{\mathcal{K}_{2}\left(z_{I}\right)}\;,\qquad W_{I}\left(z\right)=\frac{1}{4}K_{I}\sqrt{x_{I}}\;\mathcal{K}_{1}\left(z_{I}\right)z^{3}_{I}\;,$ (24)
+with $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ being the modified Bessel functions of the second kind and
+$x_{I}\equiv\frac{M^{2}_{I}}{M^{2}_{1}}\;,\qquad z_{I}\equiv z\sqrt{x_{I}}\;,\qquad K_{I}\equiv\frac{\Gamma_{I}}{H\left(T=M_{1}\right)}\;.$ (25)
+Moreover, the Hubble expansion rate $H$ is given by
+$H=\sqrt{\frac{4\pi^{3}g_{*}}{45}}\frac{M^{2}_{1}}{z^{2}M_{\rm Pl}}\;,$ (26)
+in which $M_{\rm Pl}=1.22\times 10^{19}$GeV is the Planck mass and $g_{*}\simeq 106.75$ or 228.75 (in the SM and MSSM frameworks, respectively) is the number of relativistic degrees of freedom in the early Universe.
+Last but not least, $\Lambda_{\alpha}$ is the self-energy of $\alpha$-flavored leptons, and its imaginary part is given by
+${\rm Im}\left(\Lambda_{\alpha}\right)\simeq 8\times 10^{-3}y^{2}_{\alpha}T\;,$ (27)
+with $y_{\alpha}$ being the charged-lepton Yukawa coupling.
+2.2 Some formulas about RGE effects
+As mentioned in the introduction section, the flavor-symmetry models which contain three degenerate right-handed neutrinos serve as a unique basis for realizations of the tri-resonant leptogenesis scenario after the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent.
+And the RGE effects between the flavor-symmetry and leptogenesis scales can play the role of generating the desired right-handed neutrino mass splitings.
+In the literature, most of the flavor-symmetry models have been formulated in the MSSM framework for the following reason: in order to break the flavor symmetry in a proper way, one needs to introduce some flavon fields which transform as multiplets of the flavor symmetry and develop particular VEV alignments; and the most popular and perhaps natural approach to derive the desired flavon VEV alignments is provided by the so-called F-term alignment mechanism which is realized in a supersymmetric setup [5].
+In particular, the modular flavor-symmetry models have been formulated in the MSSM framework since the modular flavor symmetry needs supersymmetry to preserve the holomorphicity of the modular form [7].
+For these reasons, we will perform our study in the MSSM framework.
+In the MSSM framework, the relevant renormalization-group equations have been given in Ref.
 [18].
-Note that at the present epoch one has the relation $s\simeq 7.04n_{\gamma}$ for the entropy and photon number densities, and correspondingly $Y_{\rm L}\simeq\eta_{\rm L}/7.04$ (with $\eta_{\rm L}=\eta_{L_{e}}+\eta_{L_{\mu}}+\eta_{L_{\tau}}$).
-2.2 Some basics about RGE effects
-As mentioned in the introduction section, the flavor-symmetry models which contain three degenerate right-handed neutrinos serve as a unique basis for realizations of the tri-resonant leptogenesis scenario after the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent.
-And the renormalization group evolution effects from high to low energies can play the role of generating the desired right-handed neutrino mass splitings.
-In the literature, most of the flavor-symmetry models have been formulated in the MSSM framework for the following reason: in order to break the flavor symmetry in a proper way, one needs to introduce some flavon fields which transform as multiplets of the flavor symmetry and develop particular VEV alignments; and the most popular and perhaps natural approach to derive the desired flavon VEV alignments is provided by the so-called F-term alignment mechanism which is realized in a supersymmetric setup [6].
-In particular, all of the modular flavor-symmetry models have been formulated in the MSSM framework since the modular flavor symmetry needs supersymmetry to preserve the holomorphicity of the modular form [7].
-For these reasons, we will perform our study in the MSSM framework.
-In the MSSM framework, the one-loop RGE evolution of the right-handed neutrino mass matrix and the Dirac neutrino Yukawa matrix $Y$ in the basis where $M_{\rm R}$ is diagonal are given by [19]
+First of all, the lepton fields can be transformed by a unitary transformation to the basis in which the right-handed neutrino mass matrix is diagonal.
+In this basis, the RGE evolution of three right-handed neutrino masses $M_{I}$ and the neutrino Yukawa coupling matrix $Y$ can be obtained by using the methods outlined in [14, 19]
 $4\pi^{2}\frac{{\rm d}M_{I}}{{\rm d}t}=\left(YY^{\dagger}\right)_{II}M_{I}\;,$
@@
 where $t=\ln\left(M_{0}/\mu\right)$ has been defined, with $M_{0}$ and $\mu$ denoting the right-handed neutrino mass scale and renormalization scale, respectively.
-$\alpha=3g^{2}_{2}+\frac{3}{5}g^{2}_{1}-{\rm Tr}\left(3Y^{\dagger}_{\rm U}Y_{\rm U}+Y^{\dagger}Y\right)$, $g_{2}$ and $g_{1}$ are the $\rm SU\left(2\right)_{\rm L}$ and $\rm U\left(1\right)_{\rm Y}$ gauge coupling constants, and $Y_{{\rm U},l}$ are the Yukawa matrices for up-type quarks and charged leptons.
-$T$ is a $3\times 3$ anti-Hermitian matrix and its entries are given by
-$T_{II}=0\;,$
+And one has $\alpha=3g^{2}_{2}+\frac{3}{5}g^{2}_{1}-{\rm Tr}\left(3Y^{\dagger}_{u}Y_{u}+Y^{\dagger}Y\right)$, where $g_{2}$ and $g_{1}$ are the $\rm SU\left(2\right)_{\rm L}$ and $\rm U\left(1\right)_{\rm Y}$ gauge coupling constants, and $Y_{u,l}$ are the Yukawa coupling matrices for the up-type quarks and charged leptons.
+$T$ is a $3\times 3$ anti-Hermitian matrix and its non-zero entries are given by
 $T_{IJ}=-T^{*}_{JI}=\frac{M_{J}+M_{I}}{M_{J}-M_{I}}{\rm Re}[(YY^{\dagger})_{IJ}]+{\rm i}\frac{M_{J}-M_{I}}{M_{J}+M_{I}}{\rm Im}[(YY^{\dagger})_{IJ}]\;,\qquad{\rm for}\;\;I\neq J\;.$ (29)
-$T_{IJ}$ is singular at the scale where $M_{J}=M_{I}$ holds unless ${\rm Re}[(YY^{\dagger})_{IJ}]=0$. To achieve this condition, we perform an additional rotation in the above basis:
-$M^{\prime}_{\rm R}=O^{\dagger}M_{\rm R}O^{*}=M_{\rm R}\;,\qquad Y^{\prime}=O^{\dagger}Y\;,$ (22)
-where $O$ is a real orthogonal matrix which can be parameterized by three angles $\theta_{1,2,3}$.
-This rotation does not affect the right-handed neutrino mass matrix due to their exact mass degeneracy.
-$YY^{\dagger}$ becomes
-$Y^{\prime}Y^{\prime\dagger}=O^{\dagger}YY^{\dagger}O\;.$ (23)
-Apparently, the three condition ${\rm Re}[(YY^{\dagger})_{IJ}]=0$ can be always satisfied by choosing suitable vuales of $\theta_{1,2,3}$.
-The RGE evolution of the right-handed neutrino mass matrix in the new basis is
-$4\pi^{2}\frac{{\rm d}M_{I}}{{\rm d}t}=\left(Y^{\prime}Y^{\prime\dagger}\right)_{II}M_{I}\;.$ (24)
-For illustration, taking account of the RGE effects from the flavor-symmetry scale $\Lambda_{\rm FS}$ to the right-handed neutrino mass scale $M_{0}$, one can obtain the following relation for the right-handed neutrino masses
-$M_{I}\left(M_{0}\right)\simeq M_{I}\left(\Lambda_{\rm FS}\right)\left[1-\frac{1}{4\pi^{2}}\left(Y^{\prime}Y^{\prime\dagger}\right)_{II}\ln\left(\frac{\Lambda_{\rm FS}}{M_{0}}\right)\right]\;,$ (25)
-yielding the following mass splittings for the right-handed neutrinos
-$M_{I}\left(M_{0}\right)-M_{J}\left(M_{0}\right)\simeq\frac{M_{I}\left(\Lambda_{\rm FS}\right)}{4\pi^{2}}\left[\left(Y^{\prime}Y^{\prime\dagger}\right)_{JJ}-\left(Y^{\prime}Y^{\prime\dagger}\right)_{II}\right]\ln\left(\frac{\Lambda_{\rm FS}}{M_{0}}\right)\;.$ (26)
-Such mass splittings among the right-handed neutrinos are comparable to their decay widths
-$\Gamma_{I}=\frac{\left(YY^{\dagger}\right)_{II}}{8\pi}M_{I}\left(M_{0}\right)\;,$ (27)
-fulfilling the condition for the resonant leptogenesis to proceed.
-Note that in obtaining Eq.
-(25), as a rough estimate, we have only considered the diagonal-term [i.e., $(YY^{\dagger})_{II}$-related] contributions to the right-handed neutrino masses.
-But the off-diagonal-term [i.e., $(YY^{\dagger})_{IJ}$-related] contributions may be equally important in the case of the right-handed neutrino masses being degenerate, so in the following calculations we will also consider them.
+It is immediate to see that $T_{IJ}$ is singular at the scale where one has $M_{J}=M_{I}$, unless ${\rm Re}[(YY^{\dagger})_{IJ}]=0$ holds.
+Fortunately, due to the exact mass degeneracy of three right-handed neutrinos, one has a freedom of rotation $N\to ON$ with $O^{T}O=I$ for them, which does not affect the unity right-handed neutrino mass matrix, but rotates the Yukawa coupling matrix to the following form
+$Y\to Y^{\prime}=OY\;.$ (30)
+Such a freedom can be utilized to assure that Eq. (28) is non-singular [20], namely,
+${\rm Re}\left[Y^{\prime}Y^{\prime\dagger}\right]={\rm Re}\left[O^{\dagger}YY^{\dagger}O\right]=0\;.$ (31)
+3 Study for a model realizing the TM1 mixing
+In this section, we perform the study for a neutrino flavor-symmetry model realizing the popular TM1 mixing [21].
+To be specific, we will consider the concrete model presented in section 3 of Ref.
+[22], which is constructed based on the ${\rm S}_{4}$ flavor symmetry in the MSSM framework.
+${\rm S}_{4}$ is the permutation group of 4 objects with 24 elements, and it has two singlet representations $\bf 1$} and $\bf 1^{\prime}$}, one two-dimensional representation $\bf 2$} and two three-dimensional representations $\bf 3$} and $\bf 3^{\prime}$}.
+In this model both the three left-handed lepton doublets and three right-handed neutrinos constitute a $\bf 3^{\prime}$} representation of ${\rm S}_{4}$, while three right-handed charged leptons are trivially $\bf 1$} representations of ${\rm S}_{4}$.
+In order to be able to yield certain predictions for the CP phases at the same time, the flavor symmetry is extended in such a way that a generalized CP symmetry [26] is also included.
+As is commonly done in the paradigm of flavor symmetries, the full symmetries including both the flavor and CP symmetries are broken (by the flavon VEVs) into different subgroups in the neutrino and charged-lepton sectors, and the mismatch between them gives rise to certain predictions for the lepton mixing angles and CP phases.
+In the model under consideration, thanks to the residual flavor and CP symmetries, the charged-lepton mass matrix is constrained to be diagonal, and the Dirac neutrino mass matrix is constrained into the following form
+$M_{\rm D}=\frac{y_{1}v_{u}v_{\xi}}{\Lambda_{\rm FS}}\left[\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix}+x\begin{pmatrix}0&1&1\\ 1&1&0\\ 1&0&1\end{pmatrix}+{\rm i}y\begin{pmatrix}0&1&-1\\ -1&0&-2\\ 1&2&0\end{pmatrix}\right$]\;, (32)
+while the right-handed neutrino mass matrix takes the same form as in Eq.
+(15).
+Here $v_{u}$ and $v_{\xi}$ are the VEVs of the up-type Higgs and flavon fields, $y_{1}$, $x$ and $y$ are real parameters.
+Given the above forms of $M_{\rm D}$ and $M_{\rm R}$, the seesaw formula yields a light neutrino mass matrix as
+$M_{\nu}=a\left(\begin{matrix}2&-1&-1\cr-1&2&-1\cr-1&-1&2\end{matrix}\right)+b\left(\begin{matrix}1&0&0\cr 0&0&1\cr 0&1&0\end{matrix}\right)+c\left(\begin{matrix}0&1&1\cr 1&1&0\cr 1&0&1\end{matrix}\right)+{\rm i}d\left(\begin{matrix}0&1&-1\cr 1&2&0\cr-1&0&-2\end{matrix}\right)\;,$ (33)
+with
+$a=m_{0}y^{2}\;,\hskip 28.45274ptb=-m_{0}(1+2x^{2}+4y^{2})\;,$
+$c=-m_{0}(2x+x^{2}+y^{2})\;,\hskip 28.45274ptd=-3m_{0}xy\;,$ (34)
+and $m_{0}=y^{2}_{1}v^{2}_{u}v^{2}_{\xi}/(M_{0}\Lambda^{2}_{\rm FS})$.
+It is easy to verify that such an $M_{\nu}$ not only keeps invariant under the TM1 transformation but also under the $\mu$-$\tau$ reflection transformation [27]:
+$M_{\nu}=R^{T}_{\rm TM1}M_{\nu}R_{\rm TM1}\hskip 14.22636pt{\rm with}\hskip 14.22636ptR_{\rm TM1}=-\frac{1}{3}\left(\begin{array}[]{ccc}1&2&2\cr 2&-2&1\cr 2&1&-2\cr\end{array}\right)\;,$ (38)
+$M_{\nu}=R^{T}_{\mu\tau}M^{*}_{\nu}R_{\mu\tau}\hskip 14.22636pt{\rm with}\hskip 14.22636ptR_{\mu\tau}=\left(\begin{array}[]{ccc}1&0&0\cr 0&0&1\cr 0&1&0\cr\end{array}\right)\;.$ (42)
+Correspondingly, the resulting neutrino mixing not only has the features of the TM1 mixing but also the features of the $\mu$-$\tau$ reflection symmetry:
+$\theta_{23}=\frac{\pi}{4}\;,\hskip 28.45274pt\delta=\pm\frac{\pi}{2}\;,\hskip 28.45274pt\alpha_{21}=0\ {\rm or}\ \frac{\pi}{2}\;,\hskip 28.45274pt\alpha_{31}=0\ {\rm or}\ \frac{\pi}{2}\;.$ (43)
+Given that the considered model is subject to further constraints on top of the TM1 mixing, if it can successfully reproduce the observed value of $Y_{\rm B}$ with the help of the RGE effects, then the models that are only subject to the TM1 mixing necessarily can also do so.
+It is found that this model can only be phenomenologically viable in the NO case and the best-fit values of $x$ and $y$ are x=0.050x=-0.050 and $y=\pm 0.233$ (corresponding to $\delta=\pm\pi/2$) [22].
+And the parameter combination $y_{1}v_{u}v_{\xi}/\Lambda_{\rm FS}$ in $M_{\rm D}$ can be expressed as $\sqrt{m_{ee}M_{0}/(1+2x^{2}+2y^{2})}$ with $m_{ee}\simeq 57.9$ meV (i.e., the effective neutrino mass for neutrinoless double beta decays).
+The most stringent bounds on neutrinoless double beta decay lifetimes come from the KamLAND-Zen [23] and GERDA [24] experiments, which constrain the effective Majorana mass as $m_{ee}<$ (28-122)meV and $m_{ee}<$ (79-180) meV, respectively.
+The model also predicts the masses of the three neutrinos: $m_{1}=57.284$ meV, $m_{2}=57.930$ meV and $m_{3}=75.488$ meV.
+For comparison, the latest Planck results constrain the sum of neutrino masses to $\sum_{i}m_{i}<$ (0.12-0.60)eV [25].
+Furthermore, the Yukawa coupling matrix can be obtained as $Y=M_{\rm D}/v_{u}$ with $v_{u}=v\tan\beta/\sqrt{1+\tan^{2}\beta}$.
+Here $\tan\beta$ is the ratio of the VEV of the up-type Higgs field $H_{u}$ to that of the down-type Higgs field $H_{d}$.
+Note that in the MSSM a moderately large $\tan\beta$ value (e.g., 30) is favored, for which the relation $v_{u}\simeq v$ holds.
+Now, we are ready to study the RGE induced tri-resonant leptogenesis for the above model: at the flavor-symmetry scale $\Lambda_{\rm FS}$, three right-handed neutrinos are completely degenerate, prohibiting leptogenesis to proceed; if there is a considerable gap between $\Lambda_{\rm FS}$ and the leptogenesis scale (roughly $M_{0}$), then the RGE effects from the former scale to the latter scale will lift the degeneracy among three right-handed neutrino masses, inducing a tri-resonant leptogenesis to work.
+Given that most parameters of the model have been fixed as described above, there are only three free parameters relevant to our study: $M_{0}$, $\Lambda_{\rm FS}$ and $\tan\beta$.
+Considering that the RGE effects only depend on $\Lambda_{\rm FS}$ in a logarithmic manner (and we have numerically checked that the dependence of the final results on the concrete value of $\Lambda_{\rm FS}$ is weak), in the following numerical calculations we will take $\Lambda_{\rm FS}=10^{16}$ GeV (i.e., around the grand unification scale) as a benchmark value.
+Now that $M_{0}$ and $\tan\beta$ are the remaining free parameters (note that the value of $\tan\beta$ will directly affect the sizes of the charged-lepton Yukawa coupling and consequently the flavor effects), in Figure 1 we have shown the RGE-induced mass splittings of the three right-handed neutrinos as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ and 50.
+The results show that the ratios $\Delta M_{IJ}/M_{0}$ (for $\Delta M_{IJ}\equiv M_{I}-M_{J}$) are nearly linearly proportional to $M_{0}$ and insensitive to $\tan\beta$.
+This is because $\Delta M_{IJ}/M_{0}$ are proportional to $YY^{\dagger}$ (see Eq.
+(28)) which in turn is proportional to $M_{0}$.
+And $YY^{\dagger}$ is almost independent of $\tan\beta$ when it takes values from 10 to 50 (as mentioned in the end of last paragraph).
+Figure 1: For the TM1 model in section 3, the mass splittings among the three right-handed neutrinos generated from the RGE effects, as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ and 50.
+Figure 2: For the TM1 model in section 3, the resulting CP asymmetries for the decays of $N_{I}$ as functions of $M_{0}$ and $\tan\beta$.
+Figure 2 has shown the resulting CP asymmetries $\varepsilon_{I}={\rm Tr}\;\varepsilon^{\left(I\right)}_{\alpha\beta}$ for the decays of $N_{I}$ as functions of $M_{0}$ and $\tan\beta$.
+The results show that $\varepsilon_{I}$ and $\varepsilon_{T}=\sum_{I}\varepsilon_{I}$ decrease with the increasing of $M_{0}$, and increase with $\tan\beta$.
+And the sign of $\varepsilon_{3}$ changes from positive to negative at $M_{0}\simeq 10^{12}$ GeV.
+Please keep in mind that here and in the following, for negative results, we have used the dashed lines to show their absolute values.
+Finally, Figure 3 has shown the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$, 20, 30 and 50.
+One can see that for $M_{0}>10^{12}$ GeV the four curves nearly converge into a single one.
+This means that in such a mass range the resulting $\eta_{B}$ is almost independent of $\tan\beta$.
+And the observed value of $\eta_{\rm B}$ can be successfully reproduced for $M_{0}\sim 7\times 10^{12}$ GeV.
+Figure 3: For the TM1 model in section 3, the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for some benchmark values of $\tan\beta$.
+The horizontal gray dashed line shows the observed value of $\eta_{\rm B}$.
 4 Study for modular ${\rm A}_{4}$ symmetry model
@@
 (15).
-On the other hand, in the case that three right-handed charged leptons are successively assigned to the singlet representations ${\bf 1}$, ${\bf 1^{\prime\prime}}$ and ${\bf 1^{\prime}}$ of $A_{4}$, the charged-lepton mass matrix takes a form as
+On the other hand, in the case that three right-handed charged leptons are successively assigned to the singlet representations $\bf 1$}, $\bf 1^{\prime\prime}$} and $\bf 1^{\prime}$} of ${\rm A}_{4}$, the charged-lepton mass matrix takes a form as
 $M_{l}=v_{d}\begin{pmatrix}\alpha Y_{1}&\alpha Y_{3}&\alpha Y_{2}\\ \beta Y_{2}&\beta Y_{1}&\beta Y_{3}\\ \gamma Y_{3}&\gamma Y_{2}&\gamma Y_{1}\end{pmatrix}$. (45)
@@
 The detailed numerical calculation shows that this model can be consistent with the neutrino oscillation experimental results in both the NO and IO cases.
-And the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 2.
+And the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 1 here.
 NO IO
@@
 ${\rm Re}(\tau)$ 0.0129 $[0,0.431]$ 0.096 $[0,0.102]$
-${\rm Im}(\tau)$ 1.824 $[0.91,1.16]\cup[1.31,1.86]$ 0.987 $[0.98,1.049]\cup[1.052,1.109]$
-$\beta/\alpha$ 205.720 $[192.39,215]\cup[3054.254093.49]$ 79.472 $[59.68,86.37]\cup[892.67,1446.02]$
-$\gamma/\alpha$ 3612.07 $[192.4,215]\cup[3066.47,4092.98]$ 1232.57 $[60.97,86.34]\cup[870.16,1443.84]$
-$|g_{2}/g_{1}|$ 2.410 $[2.398,2.71]\cup[2.95,3.86]$ 2.093 $[1.038,2.453]$
-$\arg{(g_{2}/g_{1})}$ 6.267 $[0,0.49]\cup[6.23,2\pi]$ 4.715 $[1.33,1.83]\cup[4.29,5]$
+${\rm Im}(\tau)$ 1.824 $0.91,1.16]\cup[1.31,1.86]$ 0.987 $0.98,1.049]\cup[1.052,1.109]$
+$\beta/\alpha$ 205.720 $192.39,215]\cup[3054.25,4093.49]$ 79.472 $59.68,86.37]\cup[892.67,1446.02]$
+$\gamma/\alpha$ 3612.07 $192.4,215]\cup[3066.47,4092.98]$ 1232.57 $60.97,86.34]\cup[870.16,1443.84]$
+$|g_{2}/g_{1}|$ 2.410 $2.398,2.71]\cup[2.95,3.86]$ 2.093 $[1.038,2.453]$
+$\arg{(g_{2}/g_{1})}$ 6.267 $0,0.49]\cup[6.23,2\pi]$ 4.715 $1.33,1.83]\cup[4.29,5]$
 $\alpha v_{d}$/MeV 0.5179 1.167
@@
 $\sin^{2}\theta_{23}$ 0.4638 $[0.4180,0.4676]$ 0.584 $[0.423,0.629]$
-$\delta/\pi$ 1.486 $[0.19,0.77]\cup[1.23,1.84]$ 1.458 $[0.068,1.933]$
-$\alpha_{21}/\pi$ 0.068 $[0,0.23]\cup[1.77,2]$ 0.138 $[0,0.19]\cup[1.8,2]$
+$\delta/\pi$ 1.486 $0.19,0.77]\cup[1.23,1.84]$ 1.458 $[0.068,1.933]$
+$\alpha_{21}/\pi$ 0.068 $0,0.23]\cup[1.77,2]$ 0.138 $0,0.19]\cup[1.8,2]$
 $\alpha_{31}/\pi$ 0.948 $[0.683,1.311]$ 0.997 $[0,2]$
@@
 $m_{3}$/eV 0.0661 $[0.0524,0.0716]$ 0.0013 $[0.0007,0.0015]$
+$\sum_{i}m_{i}$/eV 0.153 $[0.0991,0.1679]$ 0.1008 $[0.0942,0.1074]$
+$m_{ee}$/eV 0.0436 $[0.0202,0.0483]$ 0.0475 $[0.0439,0.0516]$
 $\chi^{2}_{\mathrm{min}}$ 30.72 $10^{-7}$
-Table 2: For the modular ${\rm A}_{4}$ symmetry model given in section 3, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in the NO and IO cases.
+Table 1: For the modular ${\rm A}_{4}$ symmetry model in section 4, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in the NO and IO cases.
 These results are taken from Tables 6 and 10 in Ref.
 [28].
-We first consider the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq.
-(10).
-In order to facilitate the leptogenesis calculation, we go to the basis where both the right-handed neutrino mass matrix and charged lepton mass matrix are diagonal with the help of $U_{\rm R}$ in Eq.
-(9) (which serves to diagonalize the the right-handed neutrino mass matrix) and another unitary matrix $U_{\rm L}$ which serves to diagonalize the charged lepton mass matrix.
-At the same time, the neutrino Yukawa coupling matrix becomes $Y^{\prime}=U^{\dagger}_{\rm R}YU_{\rm L}$ which is relevant for the leptogenesis calculations.
-With help of the Eq.
-(13) and the fitted values of the input parameters as given in Table 2, we numerically calculate the CP asymmetries in the decays of $N_{I}$ as functions of $2\mu/\Gamma_{I}$.
-The results are shown in Figure 1.
-Since $\mu$ is a free parameter, the resonance conditions $2\mu/\Gamma_{I}\simeq 1$ which enhance the CP asymmetries to their maximally allowed values can always be fulfilled.
-This will be helpful for us to reproduce the observed value of $\eta_{\rm B}$.
-Figure 2 has further shown the ratios of various collision terms in the Boltzmann equations to the Hubble constant as functions of $z=M_{0}/T$ with $M_{0}=10^{3}$ GeV and $\mu=10^{-8}$ GeV as benchmark values (see below for the reason why we have taken $\mu=10^{-8}$ GeV as a benchmark value).
-We see that the scattering collision term $\Gamma^{S}_{\rm Gauge}$ is comparable to $\Gamma^{S}_{\rm Yukawa}$, and they provide the dominant sources of the scattering collision terms in the temperature range $T>M_{0}$.
-On the other hand, in the temperature range $T<M_{0}$, the dominant term is $\Gamma^{D}_{\rm Yukawa}$ that describes the decay process.
-Then, the numerical scan finds the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ as shown in Figure 3.
-Considering that people are more interested in low-scale leptogenesis (which has the potential to be directly accessed by running or upcoming experiments), our scan over $M_{0}$ is performed in the range from $10^{3}$ GeV to $10^{6}$ GeV.
-The results show that for each value of $M_{0}$ there exist two values of $\mu$ that accommodate a successful leptogenesis, and the ratio of $\mu$ to $M_{0}$ is about $10^{-10}$ or $10^{-14}$ ($10^{-9}$ or $10^{-21}$) in the NO (IO) cases.
-For the purpose of illustration, by taking $M_{0}=10^{3}$ GeV as a benchmark value, in Figure 4 we have shown the evolutions of $\eta_{N}$ and $\eta_{B}$ as functions of $z$.
-These results are obtained by taking the initial conditions to be $\eta^{\rm ini}_{N}=\eta^{\rm ini}_{B}=0$.
-The results show that, for the benchmark value $M_{0}=10^{3}$ GeV, the observed value of $\eta_{\rm B}$ can be successfully reproduced for $\mu\simeq 10^{-8.4}$ GeV ($10^{-5.9}$ GeV) in the NO (IO) case.
-Figure 1: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq.
-(10), CP asymmetries in the decays of $N_{I}$ as functions of the ratio $2\mu/\Gamma_{I}$ in the NO (left panel) and IO (right panel) cases.
-Here $\varepsilon_{\rm tot}=\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}$ has been defined.
-Figure 2: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq.
-(10), the various collision terms in the Boltzmann equations as functions of $z=M_{0}/T$ in the NO (left panel) and IO (right panel) cases.
-These results are obtained by taking $M_{0}=10^{3}$ GeV and $\mu=10^{-8}$ GeV as benchmark values.
-Figure 3: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq.
-(10), the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ in the NO (left panel) and IO (right panel) cases.
-Figure 4: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq.
-(10), the evolutions of $\eta_{N}$ and $\eta_{B}$ as functions of $z$ for the benchmark value $M_{0}=10^{3}$ GeV and some benchmark values of $\mu$ in the NO (left panel) and IO (right panel) cases.
-Here the horizontal grey dashed line shows the observed value of $\eta_{\rm B}$ (same as below) and the vertical dotted line represents the decoupling temperature $T_{*}\simeq 131.7$ GeV of sphalerons [21].
-Then, we consider the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects.
-In this regard, Figure 5 has shown the RGE-induced mass splittings of the three right-handed neutrinos.
-In obtaining these results, we have fixed the flavor-symmetry scale $\Lambda_{\rm FS}$ at the grand unification (GUT) scale $10^{16}$ GeV.
-The results show that the ratios $\Delta M_{IJ}/M_{0}$ (for $\Delta M_{IJ}\equiv M_{I}-M_{J}$) are linearly proportional to $M_{0}$.
-This is because $\Delta M_{IJ}/M_{0}$ are proportional to $YY^{\dagger}$ [see Eq.
-(26)], which in turn is proportional to $M_{0}$ because of the seesaw relationship [see Eq.
-(6)].
-Figure 6 further shows the CP asymmetries in the decays of $N_{I}$ as functions of $M_{0}$.
-In the NO case, for $M_{0}<10^{11}$ GeV, the CP asymmetries are small and nearly independent of $M_{0}$.
-But for $M_{0}>10^{11}$ GeV, $\varepsilon_{3}$ increases rapidly with $M_{0}$, while $\varepsilon_{1}$ and $\varepsilon_{2}$ also do so after a sign reversal from being positive to negative.
-In the IO case, the CP asymmetries keep positive and change very slowly as $M_{0}$ increases.
-Finally, Figure 7 has shown the generated $\eta_{B}$ as functions of $M_{0}$.
-As one can see, the observed value of $\eta_{B}$ can be successfully reproduced for $M_{0}\simeq 6\times 10^{12}$ GeV ($4\times 10^{12}$ GeV or $6\times 10^{12}$ GeV) in the NO (IO) case.
-These results can be understood with the help of the results in Figure 6.
-Figure 5: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, $\Delta M_{IJ}/M_{0}$ (for $\Delta M_{IJ}\equiv M_{I}-M_{J}$) as functions of $M_{0}$ in the NO (left panel) and IO (right panel) cases.
-Figure 6: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the CP asymmetries in the decays of $N_{I}$ as functions of $M_{0}$ in the NO (left panel) and IO (right panel) cases.
-Figure 7: For the modular ${\rm A}_{4}$ symmetry model given in section 3, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the generated $\eta_{\rm B}$ as functions of $M_{0}$ in the NO (in the red color) and IO (in the blue color) cases.
+Now, we study the RGE induced tri-resonant leptogenesis for the above model.
+For the purpose of illustration, Figure 4 has shown the CP asymmetries for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark value of $\tan\beta=30$.
+In the NO case, for $M_{0}<10^{12}$ GeV, the CP asymmetries are nearly independent of $M_{0}$.
+But for $M_{0}>10^{12}$ GeV, the absolute value of $\varepsilon_{3}$ increases with $M_{0}$, while $\varepsilon_{1}$, $\varepsilon_{2}$ and $\varepsilon_{T}$ also do so after a sign reversal from being positive to negative.
+In the IO case, the CP asymmetries keep their signs unchanged and change very slowly as $M_{0}$ increases.
+Figure 5 has further shown the CP asymmetries for the decays of $N_{I}$ as functions of $\tan\beta$ for the benchmark value of $M_{0}=10^{12}$ GeV.
+In both the NO and IO cases, the CP asymmetries increase with $\tan\beta$.
+Figure 6 has shown the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for some benchmark values of $\tan\beta$.
+As one can see, in the NO case, the observed value of $\eta_{\rm B}$ can be successfully reproduced when $M_{0}$ is approximately between $10^{12}$ GeV and $10^{13}$ GeV.
+Figure 7 has further shown the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$.
+But in the IO case, the resulting $\eta_{\rm B}$ is always larger than the observed value, since the corresponding CP asymmetries are too large (see the right panel of Figure 4).
+Figure 4: For the modular ${\rm A}_{4}$ symmetry model in section 4, the resulting CP asymmetries for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark value of $\tan\beta=30$ in the NO (left panel) and IO (right panel) cases.
+Figure 5: For the modular ${\rm A}_{4}$ symmetry model in section 4, the resulting CP asymmetries for the decays of $N_{I}$ as functions of $\tan\beta$ for the benchmark value of $M_{0}=10^{12}$ GeV in the NO (left panel) and IO (right panel) cases.
+Figure 6: For the modular ${\rm A}_{4}$ symmetry model in section 4, the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=$10, 20, 30 and 50 in the NO (left panel) and IO (right panel) cases.
+The horizontal gray dashed line shows the observed value of $\eta_{\rm B}$.
+Figure 7: For the modular ${\rm A}_{4}$ symmetry model in section 4, in the NO case, the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$.
 5 Study for modular ${\rm S}_{4}$ symmetry model
@@
 When the three left-handed lepton doublets have a weight $k_{L}=2$ and constitute the same three-dimensional representation of ${\rm S}_{4}$ as the three right-handed neutrinos, the Dirac neutrino mass matrix will take a form as
-$M_{\rm D}=gv_{u}\left[\begin{pmatrix}0&Y_{1}&Y_{2}\\ Y_{1}&Y_{2}&0\\ Y_{2}&0&Y_{1}\end{pmatrix}+\frac{g^{\prime}}{g}\begin{pmatrix}0&Y_{5}&-Y_{4}\\ -Y_{5}&0&Y_{3}\\ Y_{4}&-Y_{3}&0\end{pmatrix}\right]\;.$ (34)
+$M_{\rm D}=gv_{u}\left[\begin{pmatrix}0&Y_{1}&Y_{2}\\ Y_{1}&Y_{2}&0\\ Y_{2}&0&Y_{1}\end{pmatrix}+\frac{g^{\prime}}{g}\begin{pmatrix}0&Y_{5}&-Y_{4}\\ -Y_{5}&0&Y_{3}\\ Y_{4}&-Y_{3}&0\end{pmatrix}\right$]\;.
+(50)
 If the three left-handed lepton doublets constitute a three-dimensional representation of ${\rm S}_{4}$ different from the three right-handed neutrinos, then the Dirac neutrino mass matrix will take a form as
-$M_{\rm D}=gv_{u}\left[\begin{pmatrix}0&-Y_{1}&Y_{2}\\ -Y_{1}&Y_{2}&0\\ Y_{2}&0&-Y_{1}\end{pmatrix}+\frac{g^{\prime}}{g}\begin{pmatrix}2Y_{3}&-Y_{5}&-Y_{4}\\ -Y_{5}&2Y_{4}&-Y_{3}\\ -Y_{4}&-Y_{3}&2Y_{5}\end{pmatrix}\right]\;.$ (35)
+$M_{\rm D}=gv_{u}\left[\begin{pmatrix}0&-Y_{1}&Y_{2}\\ -Y_{1}&Y_{2}&0\\ Y_{2}&0&-Y_{1}\end{pmatrix}+\frac{g^{\prime}}{g}\begin{pmatrix}2Y_{3}&-Y_{5}&-Y_{4}\\ -Y_{5}&2Y_{4}&-Y_{3}\\ -Y_{4}&-Y_{3}&2Y_{5}\end{pmatrix}\right$]\;.
+(51)
 On the other hand, for the minimal (in terms of weights) possibility, the charged-lepton mass matrix takes a form as
-$M_{l}=v_{d}\begin{pmatrix}\alpha Y_{3}&\alpha Y_{5}&\alpha Y_{4}\\ \beta\left(Y_{1}Y_{4}-Y_{2}Y_{5}\right)&\beta\left(Y_{1}Y_{3}-Y_{2}Y_{4}\right)&\beta\left(Y_{1}Y_{5}-Y_{2}Y_{3}\right)\\ \gamma\left(Y_{1}Y_{4}+Y_{2}Y_{5}\right)&\gamma\left(Y_{1}Y_{3}+Y_{2}Y_{4}\right)&\gamma\left(Y_{1}Y_{5}+Y_{2}Y_{3}\right)\\ \end{pmatrix}\;.$ (36)
+$M_{l}=v_{d}\begin{pmatrix}\alpha Y_{3}&\alpha Y_{5}&\alpha Y_{4}\\ \beta\left(Y_{1}Y_{4}-Y_{2}Y_{5}\right)&\beta\left(Y_{1}Y_{3}-Y_{2}Y_{4}\right)&\beta\left(Y_{1}Y_{5}-Y_{2}Y_{3}\right)\\ \gamma\left(Y_{1}Y_{4}+Y_{2}Y_{5}\right)&\gamma\left(Y_{1}Y_{3}+Y_{2}Y_{4}\right)&\gamma\left(Y_{1}Y_{5}+Y_{2}Y_{3}\right)\\ \end{pmatrix}\;$.
+(52)
 In the above three equations, $g$, $g^{\prime}$, $\alpha$, $\beta$ and $\gamma$ are simply dimensionless parameters, while $Y_{i}$ are modular forms of weight 2 and level 4 whose explicit expressions have been formulated in Ref.
@@
 The cases belonging to the first pair are denoted as Case A, while those belonging to the second pair are denoted as Case B.
-For these two cases, the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 3 [22].
+For these two cases, the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 2 here [29].
 For $M_{\rm D}$ in Eq.
@@
 In the NO case, the numerical search only finds one pair of local minima of $\Delta\chi^{2}$ corresponding to one pair of distinct values of $\tau$, and the cases belonging to this pair are denoted as Case E.
-For these three cases, the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 4 [22].
+For these three cases, the resulting predictions for the best-fit values and the allowed ranges of the input parameters and observables are collected in Table 3 here [29].
 Case A (NO) Case B (NO)
@@
 ${\rm Re}(\tau)$ $\pm 0.1045$ $\pm(0.09378-0.1128)$ $\mp 0.109$ $\mp(0.103-0.1197)$
-${\rm Im}(\tau)$ 1.01 $1.004-1.018$ 1.005 $0.9988-1.008$
-$\beta/\alpha$ 9.465 $7.693-12.39$ 0.03306 $0.02529-0.04074$
-$\gamma/\alpha$ 0.002205 $0.001941-0.002472$ 0.0001307 $0.0000982-0.0001663$
-${\rm Re}(g^{\prime}/g)$ 0.233 $-0.02544-0.4417$ 0.4097 $0.3241-0.5989$
+${\rm Im}(\tau)$ 1.01 1.0041.0181.004-1.018 1.005 0.99881.0080.9988-1.008
+$\beta/\alpha$ 9.465 7.69312.397.693-12.39 0.03306 0.025290.040740.02529-0.04074
+$\gamma/\alpha$ 0.002205 0.0019410.0024720.001941-0.002472 0.0001307 0.00009820.00016630.0000982-0.0001663
+${\rm Re}(g^{\prime}/g)$ 0.233 0.025440.4417-0.02544-0.4417 0.4097 0.32410.59890.3241-0.5989
 ${\rm Im}(g^{\prime}/g)$ $\pm 0.4924$ $\pm(-0.6046-0.5751)$ $\mp 0.5745$ $\mp(0.5436-0.5944)$
@@
 $g^{2}v_{u}^{2}/M_{0}$ [eV] 0.00467 0.004014
-$m_{e}/m_{\mu}$ 0.004802 $0.00422-0.005383$ 0.004802 $0.004211-0.005384$
-$m_{\mu}/m_{\tau}$ 0.0565 $0.04317-0.06961$ 0.05649 $0.04318-0.06962$
-$\sin^{2}\theta_{12}$ 0.305 $0.2656-0.3449$ 0.305 $0.2662-0.3455$
-$\sin^{2}\theta_{13}$ 0.02125 $0.01912-0.02383$ 0.02125 $0.01914-0.02383$
-$\sin^{2}\theta_{23}$ 0.551 $0.4838-0.5999$ 0.551 $0.4322-0.601$
-$m_{1}$ [eV] 0.01746 $0.01185-0.02143$ 0.02074 $0.01918-0.02428$
-$m_{2}$ [eV] 0.01945 $0.01473-0.02307$ 0.02244 $0.02101-0.02574$
-$m_{3}$ [eV] 0.05288 $0.05075-0.05452$ 0.05406 $0.05314-0.05577$
+$m_{e}/m_{\mu}$ 0.004802 0.004220.0053830.00422-0.005383 0.004802 0.0042110.0053840.004211-0.005384
+$m_{\mu}/m_{\tau}$ 0.0565 0.043170.069610.04317-0.06961 0.05649 0.043180.069620.04318-0.06962
+$\sin^{2}\theta_{12}$ 0.305 0.26560.34490.2656-0.3449 0.305 0.26620.34550.2662-0.3455
+$\sin^{2}\theta_{13}$ 0.02125 0.019120.023830.01912-0.02383 0.02125 0.019140.023830.01914-0.02383
+$\sin^{2}\theta_{23}$ 0.551 0.48380.59990.4838-0.5999 0.551 0.43220.6010.4322-0.601
+$m_{1}$ [eV] 0.01746 0.011850.021430.01185-0.02143 0.02074 0.019180.024280.01918-0.02428
+$m_{2}$ [eV] 0.01945 0.014730.023070.01473-0.02307 0.02244 0.021010.025740.02101-0.02574
+$m_{3}$ [eV] 0.05288 0.050750.054520.05075-0.05452 0.05406 0.053140.055770.05314-0.05577
+$\sum_{i}m_{i}$ [eV] 0.0898 $[0.07735,0.09887]$ 0.09724 $[0.0935,0.1056]$
+$m_{ee}$ [eV] 0.01699 $[0.01177,0.02002]$ 0.01983 $[0.01847,0.02275]$
 $\delta/\pi$ $\pm 1.314$ $\pm(1.249-1.961)$ $\pm 1.919$ $\pm(1.882-1.977)$
@@
 $N\sigma\equiv\sqrt{\Delta\chi^{2}}$ 0.02005 0.02435
-Table 3: For the model given in section 4 with $M_{\rm D}$ in Eq.
+Table 2: For the modular ${\rm S}_{4}$ symmetry model in section 5 with $M_{\rm D}$ in Eq.
 (50), in the phenomenologically viable NO case, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in Case A and Case B.
@@
 ${\rm Re}(\tau)$ $\mp 0.1435$ $\mp(0.1222-0.168)$ $\pm 0.179$ $\pm(0.1589-0.199)$ $\mp 0.4996$ $\mp(0.48-0.5084)$
-${\rm Im}(\tau)$ 1.523 $1.088-1.594$ 1.397 $1.236-1.529$ 1.309 $1.246-1.385$
-$\beta/\alpha$ 17.82 $9.32-23.66$ 15.35 $10.79-21.09$ 0.000243 $0.0002004-0.0002864$
-$\gamma/\alpha$ 0.003243 $0.00227-0.003733$ 0.002924 $0.002443-0.003459$ 0.03335 $0.02799-0.03926$
-${\rm Re}(g^{\prime}/g)$ $-0.8714$ $-(0.7956-1.148)$ $-1.32$ $-(1.131-1.447)$ $-0.06454$ $-(0.01697-0.1215)$
+${\rm Im}(\tau)$ 1.523 1.0881.5941.088-1.594 1.397 1.2361.5291.236-1.529 1.309 1.2461.3851.246-1.385
+$\beta/\alpha$ 17.82 9.3223.669.32-23.66 15.35 10.7921.0910.79-21.09 0.000243 0.00020040.00028640.0002004-0.0002864
+$\gamma/\alpha$ 0.003243 0.002270.0037330.00227-0.003733 0.002924 0.0024430.0034590.002443-0.003459 0.03335 0.027990.039260.02799-0.03926
+${\rm Re}(g^{\prime}/g)$ 0.8714-0.8714 (0.79561.148)-(0.7956-1.148) 1.32-1.32 (1.1311.447)-(1.131-1.447) 0.06454-0.06454 (0.016970.1215)-(0.01697-0.1215)
 ${\rm Im}(g^{\prime}/g)$ $\mp 2.094$ $\mp(1.409-2.182)$ $\pm 1.733$ $\pm(1.306-2.017)$ $\mp 0.569$ $\mp(0.4572-0.6564)$
@@
 $g^{2}v_{u}^{2}/M_{0}$ [eV] 0.004087 0.00447 0.0087
-$m_{e}/m_{\mu}$ 0.004797 $0.004215-0.005378$ 0.004786 $0.004221-0.005386$ 0.004797 $0.004393-0.005197$
-$m_{\mu}/m_{\tau}$ 0.05655 $0.04348-0.0698$ 0.0554 $0.04343-0.06968$ 0.05626 $0.04741-0.0654$
-$\sin^{2}\theta_{12}$ 0.303 $0.2657-0.3436$ 0.3031 $0.2657-0.3436$ 0.311 $0.2895-0.3375$
-$\sin^{2}\theta_{13}$ 0.02175 $0.01957-0.0242$ 0.02184 $0.01954-0.0242$ 0.02185 $0.02041-0.02351$
-$\sin^{2}\theta_{23}$ 0.5571 $0.4551-0.6026$ 0.5577 $0.5482-0.6013$ 0.4469 $0.43-0.4614$
-$m_{1}$ [eV] 0.0513 $0.04882-0.05207$ 0.05122 $0.05023-0.05212$ 0.01774 $0.01703-0.01837$
-$m_{2}$ [eV] 0.05201 $0.04958-0.05274$ 0.05193 $0.05098-0.05279$ 0.0197 $0.01906-0.02025$
-$m_{3}$ [eV] 0.01512 $0.00316-0.0163$ 0.01495 $0.01223-0.01649$ 0.05299 $0.05251-0.05346$
+$m_{e}/m_{\mu}$ 0.004797 0.0042150.0053780.004215-0.005378 0.004786 0.0042210.0053860.004221-0.005386 0.004797 0.0043930.0051970.004393-0.005197
+$m_{\mu}/m_{\tau}$ 0.05655 0.043480.06980.04348-0.0698 0.0554 0.043430.069680.04343-0.06968 0.05626 0.047410.06540.04741-0.0654
+$\sin^{2}\theta_{12}$ 0.303 0.26570.34360.2657-0.3436 0.3031 0.26570.34360.2657-0.3436 0.311 0.28950.33750.2895-0.3375
+$\sin^{2}\theta_{13}$ 0.02175 0.019570.02420.01957-0.0242 0.02184 0.019540.02420.01954-0.0242 0.02185 0.020410.023510.02041-0.02351
+$\sin^{2}\theta_{23}$ 0.5571 0.45510.60260.4551-0.6026 0.5577 0.54820.60130.5482-0.6013 0.4469 0.430.46140.43-0.4614
+$m_{1}$ [eV] 0.0513 0.048820.052070.04882-0.05207 0.05122 0.050230.052120.05023-0.05212 0.01774 0.017030.018370.01703-0.01837
+$m_{2}$ [eV] 0.05201 0.049580.052740.04958-0.05274 0.05193 0.050980.052790.05098-0.05279 0.0197 0.019060.020250.01906-0.02025
+$m_{3}$ [eV] 0.01512 0.003160.01630.00316-0.0163 0.01495 0.012230.016490.01223-0.01649 0.05299 0.052510.053460.05251-0.05346
+$\sum_{i}m_{i}$ [eV] 0.1184 0.1020.12080.102-0.1208 0.1181 0.11390.12120.1139-0.1212 0.09043 0.088740.091950.08874-0.09195
+$m_{ee}$ [eV] 0.0263 0.022880.045510.02288-0.04551 0.0310 0.025150.036770.02515-0.03677 0.006967 0.0064820.0072880.006482-0.007288
 $\delta/\pi$ $\pm 1.098$ $\pm(0.98-1.289)$ $\pm 1.384$ $\pm(1.271-1.437)$ $\pm 1.601$ $\pm(1.287-1.828)$
@@
 $N\sigma\equiv\sqrt{\Delta\chi^{2}}$ 0.0357 0.3811 2.147
-Table 4: For the model given in section 4 with $M_{\rm D}$ in Eq.
+Table 3: For the modular ${\rm S}_{4}$ symmetry model in section 5 with $M_{\rm D}$ in Eq.
 (51), in their respective phenomenologically viable neutrino mass ordering cases, the predictions for the best-fit values and the allowed ranges of the input parameters and observables in Case C, Case D and Case E.
@@
 [29].
-Now, in a similar way as in section 3, we study the consequences of these five cases for tri-resonant leptogenesis in the scenarios that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq.
-(10) and generated from the RGE effects, respectively.
-For the former scenario, the left panel of Figure 8 has shown the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $2\mu/\Gamma_{\rm tot}$.
-We see that the results in these five cases are very similar to one another, and the CP asymmetries can take their maximally allowed values around $2\mu/\Gamma_{\rm tot}\sim 1$.
-This will be helpful for us to reproduce the observed value of $\eta_{\rm B}$.
-The right panel of Figure 8 has shown the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$.
-The results for the five cases are very similar to one another, and the ratios of $\mu$ to $M_{0}$ are around $10^{-9}$ or $10^{-18}$ at $M_{0}=10^{3}$ GeV.
-For the latter scenario, the left panel of Figure 9 has shown the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $M_{0}$, while the right panel of Figure 9 has shown the generated $\eta_{\rm B}$ as functions of $M_{0}$.
-These results have also been obtained by taking $\Lambda_{\rm FS}=10^{16}$ GeV.
-The results show that only for Case E and $M_{0}\sim 10^{12}$ GeV can the observed value of $\eta_{\rm B}$ be successfully reproduced, while the the magnitudes of $\eta_{\rm B}$ are always larger than its observed value in the other four cases.
-It is interesting to recall that Case E holds in the NO case which is slightly preferred by the current experimental results compared to the IO case [2, 3].
-Figure 8: For the modular ${\rm S}_{4}$ symmetry model given in section 4, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq.
-(10), the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $2\mu/\Gamma_{\rm tot}$ (left panel), and the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ (right panel).
-Figure 9: For the modular ${\rm S}_{4}$ symmetry model given in section 4, in the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the total CP asymmetries $\varepsilon_{\rm tot}$ in the decays of $N_{I}$ as functions of $M_{0}$ (left panel), and the generated $\eta_{\rm B}$ as functions of $M_{0}$ (right panel).
+Figure 8: For the five cases of modular ${\rm S}_{4}$ symmetry model in section 5, the total CP asymmetries $\varepsilon_{T}$ for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ (left panel) and 3030 (right panel).
+Figure 9: For the five cases of modular ${\rm S}_{4}$ symmetry model in section 5, the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ (left panel) and 3030 (right panel).
+The horizontal gray dashed line shows the observed value of $\eta_{\rm B}$.
+Figure 10: For the five cases of modular ${\rm S}_{4}$ symmetry model in section 5, the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$.
+Now, we study the RGE induced tri-resonant leptogenesis for these five cases.
+Figure 8 has shown the total CP asymmetries $\varepsilon_{T}$ for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ and 30.
+For $M_{0}<10^{12}$ GeV, all of the CP asymmetries are almost independent of $M_{0}$, while they change very sharply for $M_{0}>10^{12}$ GeV.
+Figure 9 has shown the resulting $\eta_{\rm B}$ as functions of $M_{0}$ also for the benchmark values of $\tan\beta=10$ and 30.
+On can see that the observed value of $\eta_{\rm B}$ has chance to be reproduced successfully for appropriate values of $M_{0}$ and $\tan\beta$.
+This can be seen more clearly from Figure 10 which has shown the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$.
+The results show that for Case A and Case C there may exist two allowed values of $M_{0}$ for each allowed value of $\tan\beta$.
+The results show that for Case A and Case C there is an upper bound on $\tan\beta$ (about 45 and 28, respectively) and for each allowed value of $\tan\beta$ there may exist two allowed values of $M_{0}$.
+Importantly, Case C allows relatively low values of $M_{0}$, which may be reachable for future direct measurements.
+For the other cases, the allowed values of $M_{0}$ (about $10^{14}$, $10^{9}$ and $10^{13}$ GeV for Case B, Case D and Case E, respectively) are almost independent of $\tan\beta$.
 6 Study for modular ${\rm A}_{5}$ symmetry model
@@
 When the three left-handed lepton doublets have a weight $k_{L}=-2$ and constitute $\bf 3$$^{\prime}$ representation of ${\rm A}_{5}$, the Dirac neutrino mass matrix will take a form as
-$M_{\rm D}=y_{0}v_{u}\left(\begin{array}[]{ccc}\sqrt{3}Y_{1}&Y_{4}&Y_{3}\\ Y_{5}&-\sqrt{2}Y_{3}&-\sqrt{2}Y_{2}\\ Y_{2}&-\sqrt{2}Y_{5}&-\sqrt{2}Y_{4}\end{array}\right)\;.$ (40)
+$M_{\rm D}=y_{0}v_{u}\left(\begin{array}[]{ccc}\sqrt{3}Y_{1}&Y_{5}&Y_{2}\\ Y_{4}&-\sqrt{2}Y_{3}&-\sqrt{2}Y_{5}\\ Y_{3}&-\sqrt{2}Y_{2}&-\sqrt{2}Y_{4}\end{array}\right)\;.$ (56)
 On the other hand, for the minimal (in terms of weights) possibility, the charged-lepton mass matrix takes a form as
@@
 [32].
-Similar to the models in sections 3 and 4, in the present model the charged-lepton masses are essentially controlled by $\alpha v_{d}$, $\beta/\alpha$ and $\gamma/\alpha$, while the neutrino masses and mixing angles are mainly governed by $y^{2}_{0}v^{2}_{u}/M_{0}$, $\tau$, $\varphi_{2}$ and $\varphi_{3}$.
+Similar to the models in sections 4 and 5, in the present model the charged-lepton masses are essentially controlled by $\alpha v_{d}$, $\beta/\alpha$ and $\gamma/\alpha$, while the neutrino masses and mixing angles are mainly governed by $y^{2}_{0}v^{2}_{u}/M_{0}$, $\tau$, $\varphi_{2}$ and $\varphi_{3}$.
 The number of parameters is further reduced by enforcing the CP conservation which restricts all the Yukawa couplings together with $\varphi_{2}$ and $\varphi_{3}$ to be real (with the VEV of $\tau$ serving as the only source of CP violation).
-The numerical calculation shows that this model can be consistent with the neutrino oscillation experimental only in the NO case and it further predicts $m_{1}=0$.
-The resulting predictions for the best-fit values of the input parameters and observables are collected in Table 5.
-Re($\tau$) Im$(\tau)$ $\varphi_{2}$ $\varphi_{3}$ $\alpha\cos\beta\cdot 10^{3}$ $\beta/\alpha$ $\gamma/\alpha$ $y^{2}_{0}\sin^{2}\beta$/$M_{0}$ (eV-1) --0.3615 0.2412 0.04759 0.3731 3.368 1.310 --0.2413 0.0001639
-$m_{1}\cdot 10^{2}~{}{\rm eV}^{-1}$ $m_{2}\cdot 10^{2}~{}{\rm eV}^{-1}$ $m_{3}\cdot 10^{2}~{}{\rm eV}^{-1}$ $\sin^{2}\theta_{12}$ $\sin^{2}\theta_{13}$ $\sin^{2}\theta_{23}$ $\delta/\pi$ $(\alpha_{21}$-$\alpha_{31})/\pi$ $\chi^{2}_{\textrm{min}}$ 0 0.860 5.02 0.292 0.0228 0.449 1.63 1.68 11.1
-Table 5: For the model given in section 5, in the phenomenologically viable NO case, the predictions for the best-fit values of the input parameters and observables.
-These results are taken from Tables 8 in Ref.
+The numerical calculation shows that this model can be consistent with the neutrino oscillation experimental results only in the NO case and it further predicts $m_{1}=0$.
+The resulting predictions for the best-fit values of the input parameters and observables are collected in Table 4 here.
+[…]
+$m_{1}({\rm meV})$ $m_{2}({\rm meV})$ $m_{3}({\rm meV})$ $m_{ee}({\rm meV})$ $\sin^{2}\theta_{12}$ $\sin^{2}\theta_{13}$ $\sin^{2}\theta_{23}$ $\delta/\pi$ $(\alpha_{21}$-$\alpha_{31})/\pi$ $\chi^{2}_{\textrm{min}}$ 0 8.60 50.2 1.3 0.292 0.0228 0.449 1.63 1.68 11.1
+Table 4: For the modular ${\rm A}_{5}$ symmetry model in section 6, in the phenomenologically viable NO case, the predictions for the best-fit values of the input parameters and observables.
+These results are taken from Table 8 in Ref.
 [31].
-For this model, in a similar way as in sections 3 and 4, we study the consequences of it for tri-resonant leptogenesis in the scenarios that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq.
-(10) and generated from the RGE effects, respectively.
-For the former scenario, the left panel of Figure 10 has shown the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$.
-The results show that the ratio of $\mu$ to $M_{0}$ is around $10^{-8}$ or $10^{-10}$ at $M_{0}=10^{3}$ GeV.
-For the latter scenario, the right panel of Figure 10 has shown the generated $\eta_{\rm B}$ as functions of $M_{0}$.
-These results have also been obtained by taking $\Lambda_{\rm FS}=10^{16}$ GeV.
-The results show that the observed value of $\eta_{\rm B}$ can not be successfully reproduced.
-Figure 10: For the modular ${\rm A}_{5}$ symmetry model given in section 5, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq.
-(10), the values of $\mu$ that accommodate a successful leptogenesis as functions of $M_{0}$ (left panel).
-In the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the generated $\eta_{\rm B}$ as functions of $M_{0}$ (right panel).
+Figure 11: For the modular ${\rm A}_{5}$ symmetry model in section 6, the resulting CP asymmetries for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=$10 (left panel) and 30 (right panel).
+Figure 12: For the modular ${\rm A}_{5}$ symmetry model in section 6, the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=$10, 20, 30 and 50 (left panel).
+The horizontal gray dashed line shows the observed value of $\eta_{\rm B}$.
+the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$ (right panel).
+Now, we study the RGE induced tri-resonant leptogenesis for the above model.
+Figure 11 has shown the resulting CP asymmetries for the decays of $N_{I}$ as functions of $M_{0}$ for the benchmark values of $\tan\beta=10$ and 30.
+For $M_{0}<10^{12}$ GeV, all of the CP asymmetries are almost independent of $M_{0}$, while they change very sharply for $M_{0}>10^{12}$ GeV.
+The left panel of Figure 12 has shown the resulting $\eta_{\rm B}$ as functions of $M_{0}$ for some benchmark values of $\tan\beta$.
+The result shows that the resulting $\eta_{\rm B}$ is almost independent of $M_{0}$ when $M_{0}$ takes low values, and varies rapidly when $M_{0}$ becomes sufficiently large.
+Finally, the right panel of Figure 12 has shown the values of $M_{0}$ versus $\tan\beta$ that allow for a successful reproduction of the observed value of $\eta_{\rm B}$.
+The results show that the allowed values of $M_{0}$ can vary between $10^{7}$ and $10^{10}$ GeV for $\tan\beta$ from 10 to 50.
+Before proceeding to the summary section, we briefly discuss the potential constraints imposed by charged lepton violating processes on the models considered in this work.
+Among such processes, the most stringent limit stems from the $\mu\to e\gamma$ decay, as reported in Ref.
+[33]:
+${\mathrm{B}r}\left(\mu\to e\gamma\right)<3.1\times 10^{-13}\left(90\%\;{\mathrm{C.L.}}\right)\;.$ (58)
+Within the type-I seesaw framework, the branching ratio for $\mu\to e\gamma$ is given by [34]:
+${\mathrm{B}r}\left(\mu\to e\gamma\right)=\frac{2\alpha_{\rm em}}{32\pi}\left|\sum^{3}_{i=1}\left(RV\right)^{*}_{\mu i}\left(RV\right)_{ei}\left[G\left(X\right)-G\left(0\right)\right]\right|^{2}\;,$ (59)
+where $X\equiv(M_{0}/M_{W})^{2}$ with $M_{1,2,3}\simeq M_{0}$, and $G\left(X\right)$ is a loop integration function ranging in value from 4/3 to 10/3.
+Here, $V$ denotes the unitary matrix that diagonalizes the Majorana mass matrix of heavy right-handed neutrinos, and $R$ is determined by $R^{*}\simeq M_{\rm D}M^{-1}$.
+This branching ratio can be approximated as:
+${\mathrm{B}r}\left(\mu\to e\gamma\right)\simeq\frac{2\alpha_{\rm em}}{32\pi}\left(\frac{m_{\nu}}{M_{0}}\right)^{2}\;.$ (60)
+For the heavy neutrino mass range considered in this paper ($10^{3}$ to $10^{15}$ GeV), the branching ratio lies on the order of $10^{-51}$ to $10^{-27}$, which is far below the $10^{-13}$ upper limit from the latest MEG II experiment.
 7 Summary
-As we know, due to the peculiarity of the observed neutrino mixing pattern, there have been intensive model-building studies on neutrino-related physics with the employment of certain discrete non-Abelian symmetries.
-However, a shortcoming of the conventional flavor-symmetry models is that many flavon fields need to be introduced in order to achieve a desired breaking pattern of the flavor symmetry, and the scalar potential of these flavon fields needs a complicated symmetric shaping.
-In recent years, the idea of modular invariance which provides substantial simplifications to these problems has been proposed and become increasingly popular.
-However, for some typical modular symmetry neutrino models, the right-handed neutrino mass matrix takes a form as shown in Eq.
-(8) which gives three degenerate right-handed neutrino masses and consequently prohibits the leptogenesis to work.
+Due to the peculiarity of the observed neutrino mixing pattern, there have been intensive model-building studies on neutrino-related physics with the employment of certain discrete non-Abelian symmetries.
+Particularly, in recent years, the idea of modular invariance has been proposed and become increasingly popular.
+However, for some typical neutrino flavor-symmetry models, the right-handed neutrino mass matrix takes a form as shown in Eq.
+(15) which gives three exactly degenerate right-handed neutrino masses and consequently prohibits the leptogenesis mechanism to work successfully (to be specific, there would no CP asymmetries for the decays of right-handed neutrinos).
 But this kind of models can serve as a unique basis for realizations of the tri-resonant leptogenesis scenario: if the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent, then the tri-resonant leptogenesis scenario will be naturally realized.
-In this paper, we have considered two minimal ways to generate the desired small mass splittings among the three right-handed neutrinos: one way is to modify $M_{\rm R}$ to a form as shown in Eq.
-(10), while another way is to consider the renormalization-group corrections for the right-handed neutrino masses.
-To be specific, we have considered modular ${\rm A}_{4}$, ${\rm S}_{4}$ and ${\rm A}_{5}$ symmetry models that have a right-handed neutrino mass matrix as shown in Eq.
-(8).
-For these models, in the scenario that the mass splittings among the three right-handed neutrinos are realized by modifying $M_{\rm R}$ into the form as shown in Eq.
-(10), the observed value of $\eta_{\rm B}$ can always be successfully reproduced for appropriate values of $\mu$.
-These results can be easily understood from the fact that given that $\mu$ is a free parameter, the resonance conditions which enhance the CP asymmetries to their maximally allowed values can always be fulfilled (this is of course helpful for us to reproduce the observed value of $\eta_{\rm B}$).
-However, for the scenario that the mass splittings among the three right-handed neutrinos are generated from the RGE effects, the results will be highly restrictive.
-Having fixed the flavor-symmetry scale to be at the GUT scale, we reach the following conclusions for the considered representative models:
-  • For the modular ${\rm A}_{4}$ symmetry model given in section 3, the observed value of $\eta_{B}$ can be successfully reproduced for $M_{0}\simeq 6\times 10^{12}$ GeV ($4\times 10^{12}$ GeV or $6\times 10^{12}$ GeV) in the NO (IO) case.
-  • For the modular ${\rm S}_{4}$ symmetry model given in section 4, among the five cases, only for Case E (which holds in the NO case) and $M_{0}\sim 10^{12}$ GeV can the observed value of $\eta_{\rm B}$ be successfully reproduced.
-  • For the modular ${\rm A}_{5}$ symmetry model given in section 5, the observed value of $\eta_{\rm B}$ can not be successfully reproduced.
+In this paper, we have considered four representative neutrino flavor-symmetry models which contain three exactly degenerate right-handed neutrinos, including a neutrino flavor-symmetry model realizing the TM1 mixing, the modular ${\rm A}_{4}$ symmetry model, the modular ${\rm S}_{4}$ symmetry model, and the modular ${\rm A}_{5}$ symmetry model.
+For these models, we have performed a careful numerical calculation of the RGE induced mass splittings of right-handed neutrinos by employing the formulas as described in section 2.2 and subsequently the resulting baryon asymmetry by employing the formulas as described in section 2.1.
+The results show that, for appropriate values of the free parameters $M_{0}$ and $\tan\beta$, the observed value of $\eta_{\rm B}$ has chance to be successfully reproduced.
+To be specific, we reach the following conclusions for the considered representative models:
+  • For the TM1 model in section 3, the observed value of $\eta_{\rm B}$ can be successfully reproduced for $M_{0}\sim 7\times 10^{12}$ GeV.
+  • For the modular ${\rm A}_{4}$ symmetry model in section 4, in the NO case the observed value of $\eta_{\rm B}$ can be successfully reproduced when $M_{0}$ is approximately between $10^{12}$ GeV and $10^{13}$ GeV.
+While in the IO case the observed value of $\eta_{\rm B}$ can not be reproduced successfully since the magnitudes of CP asymmetries are too large.
+  • For the modular ${\rm S}_{4}$ symmetry model in section 5, for Case A and Case C there is an upper bound on $\tan\beta$ (about 45 and 28, respectively) and for each allowed value of $\tan\beta$ there may exist two allowed values of $M_{0}$.
+Importantly, Case C allows relatively low values of $M_{0}$, which may be reachable for future direct measurements.
+For the other cases, the allowed values of $M_{0}$ (about $10^{14}$, $10^{9}$ and $10^{13}$ GeV for Case B, Case D and Case E, respectively) are almost independent of $\tan\beta$.
+  • For the modular ${\rm A}_{5}$ symmetry model in section 6, the observed value of $\eta_{\rm B}$ can be successfully reproduced for $M_{0}$ in the range between $10^{7}$ and $10^{10}$ GeV for $\tan\beta$ from 10 to 50.
+For the four models discussed above, the energy scales required for successful leptogenesis are relatively high.
+This appears to contradict the common expectation that the resonant leptogenesis mechanism can lower the energy scale for successful leptogenesis to around the TeV scale.
+The reason lies in the distinct parameter constraints between general models and the ones studied here: In typical models, parameters are either unconstrained or subject to only loose restrictions, allowing resonant leptogenesis to effectively reduce the required energy scale to TeV or nearby values for successful leptogenesis.
+In contrast, the flavor structures of the models considered in this paper are tightly constrained by the imposed flavor symmetries which result in fewer free parameters (in fact, we are only left with the right-handed neutrino mass scale and $\tan\beta$ as free parameters) and thus stronger predictive power.
+This strong constraint has critical implications for the specific scenario of leptogenesis triggered by renormalization group running explored in our work: For these models, the TeV scale no longer accommodates successful leptogenesis.
+In fact, for some of them (see, e.g., Figures 9 and 12), a TeV-scale setup would lead to a lepton asymmetry larger than the observed value.
 Acknowledgments
-This work is supported in part by the National Natural Science Foundation of China under grant Nos.
-11605081, 12142507 and 12147214, the Natural Science Foundation of the Liaoning Scientific Committee under grant NO.
-2022-MS-314, and the Ph.
+Zhi-Cheng Liu acknowledges the support from the Ph.
 D.
@@
 BSJ2021048).
+Zhen-hua Zhao acknowledges the support from the National Natural Science Foundation of China under Grant No.
+12475112, Liaoning Revitalization Talents Program under Grant No.
+XLYC2403152, and the Basic Research Business Fees for Universities in Liaoning Province under Grant No.
+LJ212410165050.